A Complete Bibliography of the Journal of the Royal Statistical Society, Series A family: 1940–1949

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.02

Title word cross-reference

χ^{2} [316]. \sqrt{x} [261].

-Distribution [316].

1 [157, 229, 348, 395, 468, 540, 611]. 10th [540]. 15th [122]. 16th [230, 229, 302, 462]. 17th [171, 124]. 18th [85, 84, 275]. 1925 [567]. 1930 [320]. 1933 [145]. 1934 [211]. 1935 [63]. 1938 [193]. 1939 [318, 491]. 1941 [122]. 1942 [171, 253]. 1943 [275, 479, 379]. 1944 [349]. 1945 [396, 415]. 1946 [469]. 1947 [558, 611, 566]. 1948 [612, 611, 596]. 1949 [682]. 1960 [233].

2 [84, 302]. 20th [349, 348, 612, 558, 611].

22nd [303, 302, 415]. 24th [158, 157]. 26th [396, 395]. 27th [469, 468, 485]. 29th [682].

31st [157, 229, 302, 348, 395, 468, 540, 611]. 35 [17]. 38 [398].

41 [356]. 46 [525].

70 [305]. 7th [470].

Accident [40, 504]. Accidents [260]. Account [535]. Accounting [268]. Accounts [380, 600, 288, 380]. Activity [41]. Actuaries [596]. Actuelle [595]. Adams [211]. Additions [27, 48, 74, 96, 119, 136, 153, 167, 182, 201, 221, 238, 257, 272, 294, 310, 326, 362, 385, 410, 427, 441, 459, 482, 515, 554, 625, 657, 673].

Adelphi [84, 302]. Adolescent [397]. Adopted [488]. Adult [244]. Advanced [332, 506]. Advancement [7]. Advances [83, 464]. Advertising [422, 649, 196]. Affaires [594]. After [171, 235]. Agencies [356]. Agreements [434]. Agricole [42]. Agricultural [6, 58, 62, 172]. Agriculture [318, 121, 259, 245, 474, 280]. Aid [443]. Aitken [56]. Alan [38]. Alastair [564]. Alcohol [63]. Alec [378]. Alexander [671]. Alfred [199, 478]. Algebraic [298]. Allan [477]. Allen [174, 665]. Amaro [524]. Amended [488]. America [41]. American [565, 144, 479, 90, 284, 650, 35, 668, 318, 253]. American-Born [284]. among [40]. Analyse [42]. Analyser [561]. Analysing [192]. Analysis [473, 489, 249, 260, 521, 188, 432, 127, 191, 445, 371, 90, 430, 320, 649, 391, 542, 298, 380, 634, 37, 142, 148, 633]. Analytical [260]. André [42]. Animals [564]. Annual [600, 344, 85, 84, 158, 157, 229, 303, 302, 349, 348, 396, 395, 469, 468, 540, 612, 611, 682]. Anthony [439]. Application [636, 684, 504, 41]. Applications [375, 419, 248, 317, 471, 471, 473]. Applied [474]. Appliquée [594]. Approximate [173]. Approximation [316]. April [122]. Arbeiter [685]. Areas [330, 90, 209]. Arithmetical [368]. Arthur [69, 255, 475, 619, 449, 12, 285]. Articles [26, 47, 73, 95, 118, 135, 152, 166, 181, 200, 220, 237, 256, 271, 293, 309, 325, 337, 361, 384, 409, 426, 440, 458, 481, 499, 514, 532, 553, 570, 585, 605, 624, 639, 656, 672, 688]. Arts [85, 84, 302]. Ashley [404]. Asia [267]. Aspects [112, 145, 51, 518, 149, 213, 628, 486, 567]. Assessing [163, 370]. Assessment [299]. Assets [600]. Association [7, 35]. Associations [215]. Assurance [244]. Atlantic [297]. Attributes [5]. Audits [619]. Aurel [563]. Automobile [21]. Autoregressive [503]. aux [594]. Awards [370].

B [11, 192, 618, 479, 477, 8, 148, 63, 648]. Babington [8]. Back [29, 49, 76, 98, 184, 203, 222, 240, 328, 339, 364, 572, 587, 626, 689]. Balance [380, 19]. Balances [379]. Baldamus [565]. Balogh [617]. Banking [510, 616]. Banks [356]. Barger [318, 377, 398]. Barnard [88]. Based [373]. Bases [125]. Basic [226]. Basil [380]. Basis [517]. Be [470]. Beef [518]. Before [470, 235, 171, 275, 225]. Beginners [106]. Beginning [6]. Behaviour [319, 351]. Bell [126]. Belonging [446]. Belshaw [194]. Bennett [267]. Berkeley [666]. Bernonville [42]. Bertrand [42]. Better [494]. Between [443, 432, 246, 446, 260, 141, 518, 149]. Beveridge [228, 45, 354]. Big [529]. Binomial [264]. Biological [107]. Biology [249, 474]. Biometry [58]. Birth [109, 265]. Births [560]. bis [685]. Black [65]. Boddington [684]. Bogot [130]. Bonar [116]. Book [11, 21, 69, 192, 305, 447, 472, 473, 632, 376, 15, 20, 58, 59, 64, 91, 92, 196, 234, 251, 268, 283, 288, 355, 380, 451, 452, 526, 596, 633, 634, 668, 131, 252, 478, 489, 491, 524, 527, 579, 665, 669, 37, 494, 565, 566, 597, 618, 652, 636, 22, 61, 112, 144, 231, 250, 307, 479, 523, 651, 493, 614, 684, 249, 16, 17, 42, 44, 65, 87, 109, 129, 145, 147, 161, 175, 194, 217, 232, 235, 265, 266, 304, 306, 318, 334, 353, 358, 475, 477, 511, 545, 546, 548, 7, 8, 56, 57, 86, 106, 107, 108, 142, 143, 253, 510, 544, 36, 105, 282, 333, 507, 521, 616, 617, 619, 163, 567, 9, 10, 38, 39, 40, 62, 125, 126]. Book [127, 159, 174, 191, 264, 351, 352, 433, 434, 449, 525, 528, 581, 582, 504, 286, 422, 595, 18, 19, 45, 88, 89, 90, 529, 666, 667, 43, 212, 543, 594, 615, 685, 564, 580, 332, 506, 563, 60, 68, 111, 128, 130, 146, 148, 149, 162, 176, 177, 195, 213, 214, 215, 216, 267, 284, 287, 319, 320, 354, 356, 357, 377, 378, 379, 397, 398, 399, 400, 401, 420, 450, 492, 547, 600, 635, 649, 160, 233, 435, 448, 598, 599, 650, 683, 471, 490, 508, 522, 542, 12, 13, 14, 41, 63, 66, 67, 110, 193, 211, 285, 474, 476, 509, 648, 421]. Books [159]. Booth [406]. Borden [196]. Born [284, 320]. Boroughs [246]. Bradford [580]. Bray [380]. Bretherton [147]. Bretton [494]. Brinton [59]. Bristol [51, 250]. Britain [566, 17, 510, 148, 392, 82, 518, 147, 477, 582]. British [565, 651, 232, 567, 146, 357, 650, 443, 140, 170, 121, 372, 206, 343, 7]. Broadcasting [343]. Brown [211]. Brownlee [433]. Building [661, 44]. Burchardt [147]. Burden [232, 276]. Burn [89]. Burns [449]. Buros [159]. Business [544, 449, 529, 41, 452, 616, 127]. Business-Cycle [41].

C [485, 59, 596, 37, 266, 304, 334, 56, 253, 567, 9, 125, 264, 449, 528, 422, 564, 284, 476]. C.B [456]. C.B. [199, 324, 404]. C.B.E [383]. C.B.E. [275]. C.H. [415]. C.V.O. [383]. Cairncross [378]. Calculating [10]. Calculation [10]. Calculations [106]. Calvert [180]. Cameroons [61]. Canada [284]. Canadian [19, 320]. Capacity [467]. Capital [599, 67]. Capitalism [435, 110]. Capitalist [127]. Caradog [667]. Card [445]. Care [564]. Carl [634, 110]. Carnap [662]. Carr [266]. Carr-Saunders [266]. Carrying [467]. Cartwright [18]. Cases [188]. Cash [379]. Cattle [518]. Causes [595]. Cecil [112]. Celebrations [538]. Census [526, 630, 430, 505, 444, 124, 575, 505]. Censuses [683]. Centenary [35, 538]. Central [12, 285]. Centuries [296, 279]. Century [479, 244, 45]. Certain [188, 518]. chair [171, 122, 415, 485, 275]. Chambers [106, 40]. Changes [3, 82, 102, 450]. Chapter [509]. Charles [406]. Chart [192]. Charter [470]. Charts [299, 192]. Chemical [522]. Chi [108]. Chi-Square [108]. Children [163]. Chudson [401]. City [652, 88]. Clair [604]. Clara [603]. Clark [233]. Class [432, 206, 676, 577]. Classes [104]. Classification [663]. Clothing [577]. Clough [479]. Club [406]. Coal [664, 18]. Coal-Mining [18]. Coats [284]. Colin [233]. Collected [282]. Collection [559]. Collectiviste [42]. Colleges [487]. Collet [603]. Collins [619]. Colonial [632, 651]. Comments [315]. Commerce [684]. Commercial [287, 356, 420, 600]. Commission [52]. Committee [546, 319, 488]. Committees [420]. Commodity [434]. Community [307]. Companies [244]. Company [479]. Comparative [589, 90, 130, 356, 650]. Compared [6]. Comparison [313]. Comparisons [297]. Competition [618, 89]. Composition [193]. Computers [8]. Computing [600]. Comrie [10]. Concept [346]. Concerning [315, 278]. Conclusion [558]. Conditions [207, 435]. Conference [319, 280]. Conrad [671]. Consequences [296, 279]. Considerations [534]. Constituents [80]. Consumer [608, 356, 214]. Consumption [577]. Contemporaine [42]. Contemporary [13]. Contingency [5, 298]. Contribution [63]. Contributions [472]. Control [192, 645, 452, 434, 299, 543, 508]. Convened [558]. Copying [417]. Corporate [401, 379]. Corporation [21, 343]. Correlation [368, 432, 141]. Correlations [373, 352]. Corrigenda [205, 296, 274, 243]. Cost [104, 87, 546, 574, 148, 545, 319]. Cotton [502]. Cough [225]. Council [84, 157, 229, 302, 348, 395, 468, 488, 540, 520, 541, 611, 562, 578, 681]. Country [3]. County [246]. Cours [594]. Course [476]. Covariances [466]. Cramér [314]. Creator [110]. Credit [214, 67]. Cricket [369, 368, 389]. Crime [149]. Criminal [112, 397]. Crops [331]. Crum [447].

Cudmore [439]. Cumulative [417]. Current [25, 72, 94, 114, 133, 151, 165, 179, 198, 219, 270, 290, 322, 336, 360, 382, 403, 424, 437, 455, 496, 513, 531, 550, 569, 584, 602, 622, 638, 655, 686, 652]. Curve [521]. Curves [537]. Cycle [147, 41]. Cycles [127, 41, 449]. Cyprus [505].

D [58, 380, 524, 665, 129, 521, 89, 667, 60, 177, 267, 287, 600, 66]. D.Sc. [457]. Dahlberg [107]. Dalling [564]. Data [192, 473, 559, 346, 445]. Date [558]. Dauer [356]. Davey [216]. Davies [522]. Davis [191]. Day [435]. Deane [632]. Dearle [148]. Death [102]. Deaths [138, 123]. December [157, 229, 302, 348, 395, 468, 540, 611]. Décès [595]. Declining [11]. Defective [645]. Delinquency [130, 266]. della [523]. Demand [391, 21, 160]. Demobilization [486]. Demographic [651]. Démographie [595]. Demography [144]. Dension [406]. d’Enterprises [594]. Depreciation [600]. Depression [668, 450, 14]. Derivation [444]. Design [380]. Destiny [353]. Detailed [108]. Determination [319]. Deutschland [685]. Development [126, 534]. Developments [636, 416]. Diagrams [465]. Dictionary [86]. Difference [39, 446]. Different [446]. Diphteria [155]. Discussion [171, 275]. Disease [207]. Diseases [429]. Dispersal [155, 225]. Dispersion [330]. Displacement [306]. Distribution [392, 314, 188, 261, 52, 630, 560, 316, 173, 679, 450, 345, 367, 446]. Distributions [368, 563]. District [90]. Disturbed [344]. Division [394]. Docker [355]. Dodge [448]. Dony [252]. Dorothy [145]. Dot [465]. Doubling [448]. Drivers [40]. Dublin [478]. Dudding [192]. Dudley [581]. Due [195]. Dugé [42]. Duncan [65, 383, 471, 471]. Durbin [16]. During [450, 155, 225]. Dynamic [636]. Dynamics [21, 213].

Each [431]. Earners [629, 642]. Earnings [629, 642]. East [397]. Econometric [31]. Économi [42]. Economic [26, 47, 73, 95, 118, 135, 152, 166, 181, 200, 220, 237, 256, 271, 293, 309, 325, 337, 361, 384, 409, 426, 440, 458, 481, 499, 514, 532, 553, 570, 585, 605, 624, 639, 656, 672, 688, 196, 171, 207, 636, 129, 145, 567, 191, 351, 89, 529, 101, 518, 214, 420, 353]. Economics [11, 636, 44, 399, 233, 91, 251, 669, 217, 528, 128, 378, 635]. Économie [42]. Économique [42]. Economists [447, 665]. Economy [442, 267]. Edgerton [86]. Edna [511]. Education [357, 109]. Edward [615]. Effect [103]. Effects [196]. Efficiency [502]. Eighteenth [296, 244, 279]. Eighth [230, 229]. Eleanor [425]. Element [320]. Elementary [418, 471]. Elements [38]. Eleventh [396, 395]. Elimination [103]. Elizabeth [603, 193]. Ellinger [88]. Elliptic [141]. Elsas [235, 548]. Elsewhere [539]. Emphasis [90]. Empire [443, 651]. Employed [343]. Employment [477, 354, 377, 400, 451, 212, 399]. Ended [229, 302, 348, 395, 468, 540, 611, 157]. Ending [84, 157, 229, 302, 348, 395, 468, 540, 611]. Engineering [542]. England [491, 112, 45, 130, 149, 6, 246, 609]. English [112, 227, 678, 259]. Enquiry [266, 370]. Enterprise [597]. Epidemics [526]. Epstein [193]. Equality [466]. Equilibrium [15]. Erection [661]. Ernst [356]. Errata [389, 367]. Erratum [138, 642]. Errors [519]. Essay [217]. Essays [112, 43, 66]. Estatistica [524]. Estimate [188]. Estimating [276]. Estimation [208, 243, 189, 36]. État [42]. Étudiées [595]. Eugene [306, 508]. Europe [305, 306, 60, 176, 581]. Evan [16]. Examination [12, 285]. Examinations [299, 488]. Excess [131, 598]. Exchange [355]. Exeter [526]. Expansion [493]. Expectation [5, 669]. Expenditure [608, 246, 206, 676, 649, 286]. Experience [244, 356]. Experiment [505, 632]. Experimental [490]. Experimentation [433]. Experiments [589, 474]. Exploitation [421]. Exponential [264]. Export [517]. Exposition [58, 131]. Extension [314].

F [234, 355, 380, 636, 144, 16, 147, 7, 57, 521, 449, 43, 195, 448, 683]. F.I.A. [456]. Facit [10]. Factor [260]. Factory [139]. Family [676, 548]. Famine [525]. Farm [445]. Farmer [40]. Farms [467]. Fatality [390]. Fellows [520, 541, 562, 578]. Fenelon [22]. Fertility [444]. Fever [155]. Few [678]. Fiber [547]. Field [575, 142]. Fields [664, 591, 677, 630, 646]. Fifteenth [682]. Filadelfo [523]. Finance [140, 616]. Financial [84, 157, 229, 302, 348, 395, 468, 540, 611, 652, 556, 617, 88, 401, 420]. Financing [121, 356, 288]. Findlay [232]. Finney [521]. Fire [575]. Firm [529]. First [122, 415, 488, 485, 248, 476, 597]. Fiscal [187]. Fisher [376, 477, 497, 36]. Fit [503, 108]. Fitness [163]. Fitting [537]. Five [139]. Flexibility [451, 31]. Florence [565]. Fluctuations [352, 43, 214, 212]. Flux [199]. Food [280, 525, 176]. Foreign [421, 355]. Formation [67]. Formulae [141]. Foundations [660]. Fourier [563]. Fourteenth [612, 611]. Fowler [408]. Fraction [645]. Framework [251]. Frank [305, 633, 527]. Free [354, 400]. Frequency [314, 537, 446, 9]. Friedman [492]. Friedrich [379]. Fritz [67]. Front [1, 30, 50, 78, 100, 169, 185, 204, 224, 242, 312, 329, 341, 366, 388, 414, 484, 555, 573, 588, 607, 627, 641, 659, 675]. Fruit [331]. Full [477, 399, 354, 400]. Functions [141, 173]. Fundamental [231]. Fundamentals [471, 579, 471, 648]. Future [305, 186, 245, 357].

G [472, 252, 665, 669, 494, 22, 109, 147, 232, 353, 477, 8, 106, 333, 507, 40, 18, 332, 506, 448, 471]. G.C.B [405]. Games [351]. Garrigou [42]. Garrigou-Lagrange [42]. Gegenwart [685]. General [21, 85, 84, 158, 157, 230, 229, 303, 302, 349, 348, 396, 395, 470, 469, 468, 540, 612, 558, 611, 682, 491]. Generations [560]. Geoffrey [174]. Geometrical [369]. George [452, 63, 474, 407, 457]. Gérald [595]. Gerhard [39]. Germany [435, 421, 268]. Geschichte [685]. Gibbon [551]. Glass [60]. Glenn [90]. Gold [211]. Goodness [108, 503]. Goodwill [600]. Gopal [598]. Gottfried [214, 14]. Goulden [37]. Government [644]. Grace [109]. Grant [508]. Graphic [59]. Great [566, 17, 147, 477, 510, 582, 148, 450, 392, 82, 518]. Grebenik [250]. Greenhood [108]. Greenwood [687]. Gross [444]. Group [248, 317]. Growth [537, 90]. Guenault [18]. Guerreiro [524]. Guide [227, 192]. Gunnar [15, 107]. Gwilym [551].

H [69, 122, 196, 268, 288, 380, 131, 37, 275, 112, 231, 250, 194, 304, 318, 86, 143, 616, 9, 10, 174, 191, 422, 18, 111, 130, 284, 377, 397, 448, 598, 66]. H. [115, 485, 406]. Haberler [214, 14]. Hald [614]. Hall [85, 84]. Handbook [564]. Hanna [633]. Hans [318]. Happens [417]. Harald [314]. Harold [318, 105, 377, 398]. Harrod [636]. Harry [544, 18]. Hartkemeier [544]. Hartley [10]. Harvey [456]. Hawthorn [323]. Hawtrey [494, 353]. Hayek [43]. Head [246]. Health [644]. Heart [207]. Heilperi [91]. Held [122, 230, 229, 303, 302, 349, 348, 395, 415, 469, 468, 540, 612, 558, 611, 682, 85, 84, 158, 157, 396, 485, 462]. Henry [42, 160, 234, 407]. Herbert [44]. Hermann [266, 149, 215]. Hicks [161, 334, 286, 251, 286]. Hilferty [316]. Hill [580]. Hilton [511, 291]. Historical [127]. History [252, 527, 479, 510, 567, 89, 600, 435]. Hoel [507]. Holland [324]. Holland-Martin [324]. Home [358]. Hon [415, 485]. Hooker [323]. Horsefield [87]. Horst [450]. Hospitals [559]. Hours [629, 642]. House [302, 661]. Households [206]. Housing [235, 548]. Huber [498, 594]. Hubert [405]. Hugh [231]. Hultgren [668]. Hume [564]. Hundred [85, 84, 158, 157, 230, 229, 303, 302, 349, 348, 396, 395, 469, 468, 540, 612, 611, 682]. Hungary [20]. Hygiene [158, 157, 230, 229, 303, 349, 348, 396, 395, 469, 468, 540, 612, 558, 611, 682]. Hypotheses [226].

Ideas [226]. II [41, 33]. III [54]. Immediately [558]. Imperfect [618]. Implications [477]. Improvements [430]. Incidence [207, 65, 195, 155, 225]. Income [20, 634, 80, 65, 212, 398, 450, 193, 633, 528, 492]. Incomes [632]. Independent [5, 173, 492]. India [608, 177, 599]. Indian [62]. Industrial [375, 419, 663, 263, 248, 317, 307, 392, 52, 247, 43, 435, 609, 313, 502, 542, 433, 543]. Industries [330, 518, 377, 565, 162]. Industry [661, 190, 566, 664, 372, 18, 502, 501, 252, 357, 650, 522]. Infant [265]. Infectious [429]. Inference [660]. Information [77, 99, 154, 223, 241, 295, 340, 365, 387, 413, 461, 483, 533, 606, 674, 664, 591, 677, 630, 646]. Ingenieure [475]. Initiation [42]. Inquiry [538, 565]. Insolera [523]. Inspection [645, 592, 448, 139]. Instalment [214, 356]. Institute [596, 399]. Insurance [479]. Integral [7]. Intelligence [575]. Inter [352, 581, 3, 253]. Inter-American [253]. Inter-Regional [352]. Inter-War [3, 581]. Interest [43]. Intergovernmental [434]. Internal [619, 373]. International [64, 663, 643, 275, 618, 677, 18, 19, 128, 313, 211, 91, 477]. Interpretation [432, 284]. Interpreted [211]. Intervention [42].

Intra [432]. Introduction

[251, 544, 528, 490, 507, 378]. Introductory [112, 62]. Investigation [661]. Investigations [31]. Investment [147, 43, 41, 565]. Ioan [551]. Ireland [538, 539]. Irish [346]. Iron [534]. Irving [497]. Isles [146]. Issue [599].

J [192, 251, 288, 355, 380, 478, 112, 87, 161, 235, 334, 548, 57, 521, 616, 10, 62, 286, 18, 529, 685, 68, 130, 357, 471, 490, 13, 41, 648]. J. [180]. Jack [129]. Jacoby [616, 288]. James [526, 615, 116, 623, 408]. January [124]. Jeffreys [105]. Jenks [193]. Jennett [192]. Jenss [163]. Joan [217]. John [252, 358, 142, 143, 351, 195, 291, 84, 511]. Joint [575]. Jones [18, 667, 357]. Joseph [447, 633, 127]. Journey [307, 358]. July [612, 540, 611, 392, 462]. jun [211]. June [469, 85, 84, 158, 157, 230, 229, 303, 302, 349, 348, 396, 395, 468, 682]. Jurgen [435]. Just [92]. Juvenile [266, 130].

K.B.E [404, 456]. K.C.B. [171, 255]. Kaldor [649]. Kate [307]. Katona [452]. Kaulla [92]. Keith [87]. Kelley [579]. Kendall [8, 472, 332, 506]. Kenneth [109]. Kenney [57]. Kerrich [490]. Keynes [438]. Kingdom [63, 629, 642, 576, 79, 646]. Kirk [581]. Krisen [159]. Kuczynski [61, 651, 552, 685, 435]. Kulischer [306]. Kurtosis [301]. Kurtz [86]. Kuznets [193, 492].

L [447, 380, 669, 112, 161, 232, 304, 10, 89, 650, 508, 522]. Laboratory [394, 564]. Labour [300, 128, 148, 213, 435, 22, 421]. Lage [685]. Lagrange [42]. Laid [470]. Lamartine [434, 176]. Land [209]. Landsberg [318]. Lange [451]. Large [503, 299]. Large-Sample [503]. Laufenburger [42]. Lay [142]. Lay-Out [142]. League [420]. Leak [275, 350, 122]. Leake [600, 380, 600]. L’Économie [42].

Lee [579]. Length [678]. Léo [42]. Leon [320]. Lerner [633]. Leser [334]. Lester [684]. Levy [215]. Lewis [489]. Leybourne [109]. Liberal [615]. Liberal-Socialist [615]. Library [27, 48, 74, 96, 119, 136, 153, 167, 182, 201, 221, 238, 257, 272, 294, 310, 281, 326, 362, 385, 410, 427, 441, 459, 482, 515, 554, 625, 657, 673, 97, 168, 239, 311, 363, 411]. Liepmann [307]. Life [479, 244, 170, 307, 511]. Likelihood [55]. Lillian [193]. Limit [264]. Linder [475]. Liquor [63]. List [27, 48, 74, 96, 119, 136, 153, 167, 182, 201, 221, 238, 257, 272, 294, 310, 326, 362, 385, 410, 427, 441, 459, 482, 515, 554, 625, 657, 673]. Listener [343]. Literary [333]. Literature [466]. Liverpool [130]. Living [104, 194, 545, 574, 546]. Llewellyn [405, 406]. Lloyd [510]. Local [644, 286]. Location [565]. Logical [260]. London [88, 158, 157, 230, 229, 303, 349, 348, 396, 395, 469, 468, 540, 612, 558, 611, 682, 652]. Lord [415, 485, 134, 438, 292]. Lorimer [527]. Loss [170, 600]. Lotka [478]. Louis [478]. Lumière [595]. Lutz [379]. Lx [10].

M [472, 633, 16, 235, 265, 266, 306, 548, 8, 163, 68, 267, 284, 598, 599]. M. [115]. M.A. [323]. M.P. [425]. Machine [10]. Machlup [67]. Mackinnon [380]. Maclean [284]. made [465]. Maizels [350]. Major [687]. Majority [418]. Males [244]. Man [478]. Management [564, 542, 22]. Managerial [544]. Mance [69]. Mannheim [266, 149]. Manual [629, 642, 524]. Manufacturing [90]. Manuscripts [417]. Many [208, 243]. March [171]. Mare [34]. Margaret [12]. Marget [285]. Market [599, 391, 67]. Markets [88]. Marriage [444]. Martin [324]. Marxian [217]. Mathematical [2, 231, 5, 7, 83, 507, 125, 666, 464, 476, 283]. Mathematics [447, 394, 4, 33, 54, 56, 126, 57]. Mather [249]. Matière [42]. Matolcsy [20]. Matter [29, 49, 76, 98, 1, 30, 50, 78, 100, 184, 203, 222, 240, 169, 185, 204, 224, 242, 328, 339, 364, 312, 329, 341, 366, 388, 414, 484, 572, 587, 626, 555, 573, 588, 607, 689, 627, 641, 659, 675]. Matthias [20]. Maurice [332, 506]. Maverick [489]. Maximum [55]. May [275]. McLaughlin [90]. Meade [615]. Mean [5, 143]. Means [465, 519]. Measles [390, 225]. Measurement [632, 331, 160, 216]. Measuring [449]. Mechanism [615]. Medical [591, 107, 580]. Medicine [157, 229, 395, 468, 540, 611, 158, 230, 303, 349, 396, 469, 612, 558, 682]. Mediziner [475]. Meeting [81, 122, 230, 470, 415, 485, 557, 558, 462, 85, 84, 158, 157, 229, 303, 302, 349, 348, 396, 395, 469, 468, 540, 612, 611, 682]. Memorandum [262, 470]. Mendershausen [450]. Mentha [595]. Merchant [576]. Messrs. [350]. Meston [292]. Method [192, 376, 55, 537, 41, 373, 39]. Méthode [42]. Methoden [475]. Methodology [159]. Methods [545, 107, 19, 430, 331, 276, 343, 683, 522, 474, 37, 163]. Metoder [614]. Michael [91, 400]. Michel [594, 498]. Middle [244, 206]. Middle-Class [206]. Migone [253]. Migration [146]. Milk [518]. Milton [492]. Minimum [64]. Mining [664, 18, 377]. Mints [510]. Miss [425]. Mitchell [604, 449]. Model [10]. Modern [177, 10, 287]. Modernization [493]. Molina [264]. Moments [537]. Monetary [91, 275, 12, 13, 66, 285, 15]. Money [478, 234]. Monsoon [267]. Moore [174]. Morbidity [559]. Morgenstern [351]. Mortality [244, 265]. Mossé [42]. Motor [566, 40]. Motors [21]. Movements [145, 259, 60]. Mr. [122, 275, 374]. Mulky [599]. Municipal [619]. Murray [358]. Mutual [443, 479]. Mutually [5]. Myrdal [15].

N [62, 564, 148]. Nation [63]. National [632, 394, 20, 634, 663, 80, 445, 193]. Nations [431, 280, 420]. Nature [664, 591, 677, 630, 646]. Naturwissenschafter [475]. Navy [486]. Negative [432]. Neil [196, 288, 616]. Nest [34]. Net [346, 444]. Neumann [351]. Newsholme [255]. Nicholas [649]. Nine [661]. Nineteenth [45, 296, 279]. Ninth [303, 302]. Nixon [511]. No [8]. Nogaro [42]. Non [661]. Non-Traditional [661]. Normal [466]. Norwood [397]. Note [6, 300, 301, 662, 315, 189, 374, 316]. Notes [25, 72, 94, 24, 46, 71, 93, 114, 133, 151, 165, 113, 132, 150, 164, 179, 198, 219, 178, 197, 218, 236, 270, 290, 254, 269, 289, 308, 322, 336, 360, 321, 335, 359, 382, 403, 381, 402, 424, 437, 455, 423, 436, 454, 480, 496, 513, 531, 550, 495, 512, 530, 549, 569, 584, 602, 622, 568, 583, 601, 621, 638, 655, 686, 637, 654, 670]. Notestein [305]. Notices [520, 541, 562, 578]. November [558, 470, 415, 485]. Number [431]. Numbers [344, 8].

O [10, 522]. O.B.E. [180, 623]. Observations [124]. October [629, 642]. Offenders [266]. Office [128]. Official [262]. One [85, 84, 158, 157, 230, 229, 303, 302, 349, 348, 396, 395, 469, 468, 540, 612, 611, 682]. Operating [356]. Ordinary [122, 415, 485, 558]. Organization [575, 617]. Osborne [69]. Oscar [451]. Oscillation [103]. Oscillatory [472, 259, 371]. Oskar [159, 351]. Other [23, 70, 453, 620, 653, 43]. Our [172]. Outbreak [32]. Outdoor [227]. Outlay [398]. Output [346, 377, 313]. Oversea [646]. Overseas [79]. Oxford [399].

P [192, 485, 380, 565, 112, 163, 62, 352, 434, 18, 176, 397, 600]. P.C [415]. P.M [349, 396, 469]. Paper [374]. Papers [282]. Parameters [208, 243]. Part [435]. Paterson [58]. Pattern [401]. Paul [507, 543]. Pay [16]. Payments [19]. Peace [140]. Peach [543]. Pearl [117]. Pechman [633]. Pelle [544]. Penal [112]. Percentage [592]. Percy [404, 456, 397]. Period [545]. Periodic [344]. Periodical [28, 75, 120, 137, 205, 183, 202, 274, 258, 273, 327, 338, 386, 412, 428, 460, 500, 516, 571, 586, 640, 658]. Periodicals [97, 26, 47, 73, 95, 168, 118, 135, 152, 166, 239, 181, 200, 220, 237, 311, 256, 271, 293, 309, 363, 325, 337, 361, 411, 384, 409, 426, 440, 458, 481, 499, 514, 532, 553, 570, 585, 605, 624, 639, 656, 672, 688]. Peters [125]. Petition [470]. Petroleum [162]. Phyllis [632]. Physical [394, 163]. Pickard [526]. Pigou [567, 528]. Pittsburgh [90]. Plan [534]. Planning [18, 615]. Plans [275]. Plant [565]. Plummer [9]. Poisson [261, 446, 264]. Polanyi [400]. Policies [60, 146]. Policy [582, 177, 319, 420, 636, 112, 287]. Politique [42, 42]. Population [305, 526, 300, 527, 250, 392, 145, 306, 51, 581, 246, 3, 320, 4, 33, 54, 124, 509, 52, 11, 582, 60]. Populations [466, 446]. Position [6, 2, 190, 556, 130]. Post [275, 545, 420]. Post-War [275, 420]. Potentielle [595]. Poverty [265, 175]. Practical [131]. Practice [58, 539, 492]. Pre [526, 227]. Pre-Census [526]. Pre-War [227]. Prepared [319]. Present [2, 190, 435]. Presentation [59]. Presented [157, 229, 302, 348, 395, 468, 540, 611, 84]. President [546, 122, 415, 485, 171]. Press [649]. Price [296, 615, 319, 279, 92, 68, 451, 452, 162]. Prices [53, 156, 210, 277, 347, 393, 463, 536, 610, 680, 12, 285, 194, 45]. Primary [434]. Principles [231, 634, 580]. Probability [7, 662, 173, 563, 660, 105, 490, 9, 666]. Probit [521]. Problem [334, 63, 69]. Problems [300, 652, 529, 146, 12, 285]. Procedure [58]. Procedures [125]. Proceedings [81, 85, 122, 158, 230, 303, 349, 396, 415, 469, 485, 557, 612, 558, 593, 613, 682, 631, 647, 462, 666]. Process [127]. Processes [504]. Producing [518]. Production [263, 42, 352, 247, 630, 176, 522, 542]. Productivity [609, 650, 377]. Products [434]. Professional [492]. Professor [291]. Profit [600]. Profits [131, 598, 43]. Progress [175]. Progression [369]. Projections [305]. Proneness [40]. Proof [108]. Proposal [470]. Proposed [535]. Prospect [101]. Prospects [527]. Prosperity [668, 14]. Provisions [296, 279]. Public [140, 147]. Publications [23, 70, 453, 620, 653]. Punched [445]. Punched-Card [445].

Quality [192, 543, 508]. Quotient [314].

R [376, 251, 596, 665, 494, 636, 61, 651, 147, 161, 175, 334, 353, 36, 125, 286, 146, 177, 284, 287, 13, 648]. Rachel [163]. Radice [17]. Radio [422]. Radzinowicz [112]. Rae [457]. Rail [69]. Railway [177]. Railways [177]. Random [8, 504]. Rangnekar [618]. Ranking [189]. Ransom [526]. Rate [390, 109]. Rates [177, 102, 444]. Rathbone [425]. Rating [334]. Raul [253]. Raymond [288, 616, 117]. Re [211, 285]. Re-Examination [285]. Re-Interpreted [211]. Real [87]. Realistic [565]. Rearing [518]. Reasoning [38]. Receivable [288]. Received [97, 168, 239, 311, 363, 411]. Reconstitution [375]. Reconstruction [171]. Reconversion [493]. Reddaway [11]. Reference [644, 346, 445, 502, 522]. Reflections [539]. Reform [112]. Regard [172]. Reginald [216, 323]. Regional [352]. Registrar [491]. Registrar-General [491]. Regression [561]. Regulations [419]. Rejoinder [350]. Relating [664]. Relationship [260, 141, 246]. Relief [525]. Rene [552]. Report [84, 157, 229, 302, 348, 395, 468, 488, 540, 487, 611, 681, 228, 52, 420, 546]. Reports [520, 541, 562, 578]. Reproduction [444]. Research [376, 375, 419, 58, 159, 162, 343, 628, 522, 542, 575, 319]. Response [521]. Restriction [68]. Results [226, 19]. Retail [215]. Retailers [422]. Retrospect [101, 139]. Returns [28, 75, 120, 137, 205, 183, 202, 274, 258, 273, 327, 338, 386, 412, 428, 460, 500, 516, 571, 586, 640, 658]. Revenue [649]. Review [11, 21, 69, 192, 305, 447, 472, 473, 632, 376, 15, 20, 58, 59, 64, 91, 92, 196, 234, 251, 268, 283, 288, 355, 380, 451, 452, 526, 596, 633, 634, 668, 131, 252, 478, 489, 491, 524, 527, 579, 665, 669, 37, 494, 565, 566, 597, 618, 652, 636, 22, 61, 112, 144, 231, 250, 307, 479, 523, 651, 493, 614, 684, 249, 16, 17, 42, 44, 65, 87, 109, 129, 145, 147, 161, 175, 194, 217, 232, 235, 265, 266, 304, 306, 318, 334, 353, 358, 475, 477, 511, 545, 546, 548, 7, 8, 56, 57, 86, 106, 107, 108, 142, 143, 253, 510, 544, 36, 105, 282, 333, 507, 521, 616, 617, 619, 163, 567, 9, 10, 38, 39, 40, 62, 125, 126]. Review [127, 159, 174, 191, 264, 351, 352, 433, 434, 449, 525, 528, 581, 582, 504, 286, 422, 595, 18, 19, 45, 88, 89, 90, 529, 666, 667, 43, 212, 543, 594, 615, 685, 564, 580, 332, 506, 563, 60, 68, 111, 128, 130, 146, 148, 149, 162, 176, 177, 195, 213, 214, 215, 216, 267, 284, 287, 319, 320, 354, 356, 357, 377, 378, 379, 397, 398, 399, 400, 401, 420, 450, 492, 547, 600, 635, 649, 160, 233, 435, 448, 598, 599, 650, 683, 471, 490, 508, 522, 542, 12, 13, 14, 41, 63, 66, 67, 110, 193, 211, 285, 474, 476, 509, 648, 416, 421, 235]. Reviews [159]. Rheumatic [207]. Rhodes [266]. Rice [267]. Richard [265]. Road [69, 628]. Robert [61, 42, 324, 552]. Robertson [66]. Robinson [44, 217]. Rodney [649]. Romig [448]. Rostas [161, 232, 650]. Rowntree [175]. Royal [85, 84, 158, 157, 230, 229, 281, 303, 302, 349, 348, 396, 395, 469, 468, 540, 612, 558, 611, 682, 171, 275, 248, 317, 462, 52, 486]. Rt. [415, 485]. Rubber [677]. Rudimentary [447]. Rudolf [92]. Rural [330, 209].

Rurale [42]. Russell [108]. Russia [124]. Rutherford [147].

S [283, 634, 669, 618, 147, 146, 213]. Sabielny [10]. Safety [628]. Salaries [82]. Sales [195]. Salter [69]. Salvage [370]. Sam [377]. Same [558]. Sample [503]. Sampling [8, 315, 331, 278, 519, 471, 416, 417, 448, 448, 683, 509]. Sandage [422]. Sargant [565]. Sastry [62]. Saulnier [288, 616, 13]. Saunders [266]. Savings [17]. Scarborough [648]. Scarlet [155]. Scatter [465]. Schemes [645, 503]. School [158, 157, 230, 229, 303, 349, 348, 396, 395, 469, 468, 540, 612, 558, 611, 682]. Schultz [160]. Schumpeter [447, 127]. Schurr [377]. Schut [68]. Science [7]. Scientific [542]. Scores [389, 369, 368]. Scotland [535, 609]. Script [299]. Second [317, 175, 159]. Sections [375, 419, 593, 613, 631, 647]. Secular [102]. Security [297]. Sedley [439]. Seebohm [175]. Seed [131]. Selected [542]. Selection [330]. Sequential [645, 473]. Series [489, 103, 371, 344, 373, 472, 174, 191]. Service [644]. Session [122, 415, 485, 248, 317]. Sessional [157, 229, 302, 348, 395, 468, 540, 611]. Seventh [158, 157]. Sewell [380]. Shackle [669]. Shannon [250]. Sheasby [380]. Sheet [380]. Shepard [479]. Sheppard [7]. Shipping [576]. Shirras [232]. Short [227, 435]. Shoup [634]. Shown [244]. Sickness [504]. Side [431]. Sidney [633]. Sigmoid [521]. Significance [307, 143, 174, 187]. Silverman [649]. Simmonds [596]. Simon [492, 193]. Simple [58]. Since [244]. Singer [268, 111]. Single [448]. Sir [199, 171, 405, 255, 551, 383, 404, 456]. Situation [32]. Sixth [85, 84]. Size [565, 246, 529]. Skew [368]. Small [529]. Smith [405, 471, 143, 406, 8]. Smoothing [489]. Snedecor [474]. Snyder [110]. Social [296, 538, 251, 297, 342, 207, 175, 279, 145, 667, 149]. Socialist [615]. Society [85, 84, 302, 538, 85, 84, 158, 157, 230, 229, 281, 303, 302, 349, 348, 396, 395, 470, 469, 468, 540, 612, 558, 611, 682, 171, 275, 248, 317, 462, 556, 354]. Solution [615]. Some [296, 297, 80, 661, 636, 31, 368, 539, 51, 226, 534, 628, 172, 124, 279, 112]. Sopravvivenza [523]. Sources [664, 591, 677, 630, 646]. Souther [163]. Soviet [305, 527, 124]. Special [470, 558, 644, 346, 664, 591, 677, 445, 630, 90, 646, 502, 522]. Specification [344]. Speculation [417]. Spensley [180]. Square [5, 108]. Stage [519]. Stages [489]. Stamp [134]. Standard [519, 211]. Standardized [268]. Standards [194, 286]. States [320, 398, 41, 510]. Statistic [101]. Statistical [473, 35, 85, 84, 26, 47, 73, 95, 24, 46, 71, 93, 158, 157, 118, 135, 152, 166, 113, 132, 150, 164, 230, 229, 181, 200, 220, 237, 178, 197, 218, 236, 281, 303, 302, 256, 271, 293, 309, 254, 269, 289, 308, 349, 348, 325, 337, 361, 321, 335, 359, 396, 395, 384, 409, 381, 402, 469, 468, 426, 440, 458, 481, 423, 436, 454, 480, 540, 499, 514, 532, 553, 495, 512, 530, 549, 612, 558, 611, 570, 585, 605, 624, 568, 583, 601]. Statistical [621, 682, 639, 656, 672, 688, 637, 654, 670, 171, 491, 37, 275, 248, 317, 664, 34, 462, 429, 591, 142, 253, 36, 333, 521, 38, 127, 159, 677, 535, 630, 646, 518, 226, 148, 284, 649, 534, 628, 517, 660, 542, 41, 486, 416, 538, 376, 58, 249, 56, 86, 106, 107, 125, 508, 522, 41, 474]. Statistiche [475]. Statistician [501]. Statisticians [447]. Statistics [262, 487, 2, 644, 661, 643, 227, 678, 545, 539, 664, 591, 83, 62, 677, 186, 576, 630, 666, 646, 543, 418, 320, 464, 187, 444, 172, 471, 501, 590, 399, 283, 579, 231, 57, 544, 507, 580, 332, 506, 476, 648, 665, 684, 304, 504]. Statistique [42, 594]. Statistiques [594]. Statistiske [614]. Steel [162, 534].

Steelmaking [89]. Steindl [529]. Stephen [20]. Stock [467, 67]. Stock-Carrying [467]. Stocks [397]. Story [511]. Straw [252]. Street [84]. Structure [565, 372, 401]. Stuart [671]. Student’s [282, 107]. Studies [144, 129, 617, 128, 635]. Study [472, 390, 368, 265, 429, 333, 40, 434, 18, 19, 89, 130, 319, 377, 502, 63, 62]. Studying [373]. Subsidies [574]. Successive [560]. Suggestions [172]. Suisse [595]. Suitable [330]. Sum [173]. Sundararama [62]. Sunspot [344]. Supplemental [470]. Survey [342, 651, 175, 445, 430, 518, 64, 547]. Surveys [519, 416, 667, 683]. Susan [163]. Swaine [145]. Swedish [145]. Symbols [86]. Symposium [666]. System [445].

T [129, 617, 126, 191, 212, 564, 397]. Table [261]. Tables [645, 7, 8, 448, 298]. Tabulation [505]. Targets [517]. Tattersfield [521]. Tax [131]. Taxation [31, 161, 276, 232, 195, 598]. Taxes [65]. Teaching [487, 590]. Technical [575]. Technique [58]. Techniques [545, 542]. Tendencies [539]. Tenth [349, 348]. Teorica [523]. Terms [86]. Test [108, 174, 503]. Testing [226, 466, 41]. Tests [143]. Their [142, 373, 636, 684, 125, 504, 600]. Them [143]. Theorem [314, 315, 278]. Theoretical [127]. Theories [41]. Theory [234, 636, 539, 510, 260, 36, 662, 43, 226, 332, 506, 195, 160, 598, 471, 490, 12, 285, 12, 13, 66, 285, 92, 105, 351]. Third [535]. Thirteenth [540]. Thirty [139]. Thirty-Five [139]. Thomas [62, 145]. Thor [668]. Three [213]. Thursday [469, 558]. Time [390, 174, 191, 103, 371, 560, 679, 345, 367, 344, 373, 472, 489]. Time-Series [103, 371, 373, 472]. Times [661, 526]. Timoshenko [352]. Tin [68, 68]. Tinbergen [41, 41]. Tintner [39]. Tippett [304]. Titmuss [265]. Tiwari [177, 177, 287].

Tocher [408]. Togoland [61]. Total [444]. Town [3]. Townsend [526]. Tracts [8]. Trade [147, 79, 646, 215, 618, 400]. Traditional [661]. Training [357]. Trans [297]. Trans-Atlantic [297]. Transforms [563]. Transport [69]. Transportation [668]. Treatment [521, 600]. Trees [216]. Treloar [38]. Trend [300, 390, 103]. Trends [90]. Trials [142]. Tropical [158, 157, 230, 229, 303, 349, 396, 395, 469, 468, 540, 612, 558, 611, 682]. Troupe [380]. Truesdell [320]. Truman [579]. Tudor [357]. Tuesday [85, 158, 230, 303, 349, 396, 612]. Turner [112]. Twelfth [469, 468, 45]. Two [314, 173, 446, 519]. Two-Stage [519]. Type [344]. Types [661]. Typewriter [465].

U [161, 334, 286]. U.F.A.W [564]. U.S. [443]. U.S.S.R. [300]. Udny [333]. Unemployed [111]. Unemployment [32, 111]. Unequal [188]. Union [305, 527]. United [510, 320, 398, 41, 63, 629, 642, 576, 79, 646, 280]. Universities [487]. University [487, 399]. Use [143, 445]. Uses [661]. Utilization [209].

V [334, 352, 594, 60, 267]. Validity [589]. Valuation [334]. Value [478, 600]. Values [380]. Varga [20]. Variables [314]. Variance [188, 261, 432]. Variances [466]. Variate [39]. Variates [173, 446]. Variations [676]. Vegetable [227]. Verdier [623]. Verrijn [671]. VII [7]. Violence [138, 123]. Vocabulary [333]. Vol [435, 471]. Volume [77, 99, 154, 223, 241, 295, 340, 365, 387, 413, 461, 483, 533, 606, 674, 263, 346, 7, 247]. Voorhis [125]. Voting [418].

W [11, 192, 305, 447, 115, 268, 565, 112, 44, 7, 510, 125, 174, 564, 111, 213, 397, 12, 285, 474, 648]. W.C [84, 157, 229, 302, 348, 395, 468, 540, 611].

Wage [629, 642, 64]. Wage-Earners [629, 642]. Wages [194, 45]. Wagner [648]. Wales [491, 6, 246, 130]. Wallace [597]. Wallis [174]. Walshaw [146]. Walter [144, 401]. War [171, 275, 227, 161, 235, 545, 140, 3, 128, 148, 420, 635, 431, 581, 32, 6, 16, 87]. Warfare [129]. Wars [679, 345, 367, 170, 149]. Wartime [342]. Wasting [600]. Wealth [265, 161]. Weatherburn [476]. Wednesday [682]. Weighted [188]. Weighting [188]. Wesley [449, 604]. Western [176]. Wheat [352]. White [109]. Wholesale [53, 156, 210, 277, 347, 393, 463, 536, 610, 680]. Whooping [225]. Whooping-Cough [225]. Wickizer [267]. Wilks [283]. Willard [59]. Willcox [144]. William [199, 597, 623, 45, 354, 211]. Wilson [383, 316, 212, 63]. Winfield [112]. Wintner [563]. Wisconsin [633]. Wishart [142]. Withdrawn [470]. Within [432]. Wolfenden [231]. Wolfer [344]. Women [190]. Wood [407]. Woods [494]. Wool [442]. Woolton [415, 485]. Worcestershire [296, 279]. Worden [564]. Words [678]. Work [307, 575]. Workers [376]. Working [629, 104, 642, 676, 577]. World [442, 148, 420, 352, 547]. Worse [494]. Woytinsky [213]. Written [299].

x [261]. XXIV [8].

Yates [434, 176, 683]. Year [84, 157, 229, 302, 348, 395, 468, 540, 611]. Yearbook [253, 159]. Years [491, 139, 581]. Yield [31]. York [479, 175]. Young [397, 266]. Yule [333, 374].

References

Anonymous:1940:FMa

[1] Anonymous. Front matter. Journal of the Royal Statistical Society, 103(1): i–viii, ???? 1940. ISSN 0952-8385.

                                             URL http://www.jstor.org/stable/
                                             2980548.

Bartlett:1940:PPM

[2] M. S. Bartlett. The present position of mathematical statistics. Journal of the Royal Statistical Society, 103(1):1–29, ???? 1940. ISSN 0952-8385. URL http:

                                             //www.jstor.org/stable/2980549.

MacDougall:1940:IWP

[3] G. D. A. MacDougall. Inter-war population changes in town and country. Journal of the Royal Statistical Society, 103 (1):30–60, ???? 1940. ISSN 0952-8385.

                                             URL http://www.jstor.org/stable/
                                             2980550.

Rhodes:1940:PM

[4] E. C. Rhodes. Population mathematics — I. Journal of the Royal Statistical Society, 103(1):61–89, ???? 1940. ISSN

                                             0952-8385. URL http://www.jstor.
                                             org/stable/2980551.

Geary:1940:MEM

[5] R. C. Geary. The mathematical expectation of the mean square contingency when the attributes are mutually independent. Journal of the Royal Statistical Society, 103(1):90–91, ???? 1940. ISSN

                                             0952-8385. URL http://www.jstor.
                                             org/stable/2980552.

Anonymous:1940:NAP

[6] Anonymous. Note on the agricultural position in England and Wales, as compared with the beginning of the War in 1914. Journal of the Royal Statistical Society, 103(1):92–93, ???? 1940. ISSN

                                             0952-8385. URL http://www.jstor.
                                             org/stable/2980553.

H:1940:BRBa

[7] H. O. H. Book review: British Association for the Advancement of Science Mathematical Tables. Volume VII. The Probability Integral, by W. F. Sheppard. Journal of the Royal Statistical Society, 103(1):94–96, ???? 1940. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980554.

H:1940:BRBb

[8] H. O. H. Book review: Tracts for Computers. No. XXIV. Tables of Random Sampling Numbers, by M. G. Kendall; B. Babington Smith. Journal of the Royal Statistical Society, 103(1):96–97, ???? 1940. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980555.

K:1940:BRBa

[9] M. G. K. Book review: Probability and Frequency, by H. C. Plummer. Journal of the Royal Statistical Society, 103 (1):97, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980556.

K:1940:BRBb

[10] M. G. K. Book review: Modern Machine Calculation with the Facit Calculating Machine Model Lx, by H. Sabielny; L. J. Comrie; H. O. Hartley. Journal of the Royal Statistical Society, 103(1):97–98, ???? 1940. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980557.

A:1940:BRBa

[11] J. E. A. Book review: The Economics of a Declining Population, by W. B. Reddaway. Journal of the Royal Statistical Society, 103(1):98–99, ???? 1940. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980558.

W:1940:BRBa

[12] J. E. W. Book review: The Theory of Prices. An Examination of the Central Problems of Monetary Theory, by Arthur W. Margaret. Journal of the Royal Statistical Society, 103(1):99–101, ???? 1940. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2980559.

W:1940:BRBb

[13] J. E. W. Book review: Contemporary Monetary Theory, by R. J. Saulnier. Journal of the Royal Statistical Society, 103(1):101–102, ???? 1940. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980560.

W:1940:BRBc

[14] J. E. W. Book review: Prosperity and Depression, by Gottfried von Haberler. Journal of the Royal Statistical Society, 103(1):102, ???? 1940. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980561.

B:1940:BRBa

[15] G. A. B. Book review: Monetary Equilibrium, by Gunnar Myrdal. Journal of the Royal Statistical Society, 103(1): 103–104, ???? 1940. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980562.

G:1940:BRBa

[16] C. O. G. Book review: How to Pay for the War, by Evan F. M. Durbin. Journal of the Royal Statistical Society, 103(1): 104–105, ???? 1940. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980563.

G:1940:BRBb

[17] H. W. G. G. Book review: Savings in Great Britain, 1922–35, by E. A. Radice. Journal of the Royal Statistical Society, 103(1):105–106, ???? 1940. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980564.

M:1940:BRBa

[18] H. W. M. Book review: The Coal-Mining Industry, an International Study in Planning, by J. Harry Jones; G. Cartwright; P. H. Guenault. Journal of the Royal Statistical Society, 103(1): 106–108, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980565.

M:1940:BRBb

[19] H. W. M. Book review: The Canadian Balance of International Payments, a Study of Methods and Results. Journal of the Royal Statistical Society, 103(1): 108–111, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980566.

B:1940:BRBb

[20] G. A. B. Book review: The National Income of Hungary, by Matthias Matolcsy; Stephen Varga. Journal of the Royal Statistical Society, 103(1):111–112, ???? 1940. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980567.

A:1940:BRBb

[21] S. J. A. Book review: The Dynamics of Automobile Demand, by General Motors Corporation. Journal of the Royal Statistical Society, 103(1):112, ???? 1940. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980568.

E:1940:BRBa

[22] D. P. E. Book review: Management and Labour, by K. G. Fenelon. Journal of the Royal Statistical Society, 103(1): 112–113, ???? 1940. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980569.

Anonymous:1940:ONPa

[23] Anonymous. Other new publications. Journal of the Royal Statistical Society, 103(1):113–114, ???? 1940. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980570.

Anonymous:1940:SNa

[24] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 103(1): 115–128, ???? 1940. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980571.

Anonymous:1940:CNa

[25] Anonymous. Current notes. Journal of the Royal Statistical Society, 103(1): 129–130, ???? 1940. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980572.

Anonymous:1940:SEAa

[26] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 103(1): 131–136, ???? 1940. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980573.

Anonymous:1940:LALa

[27] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 103(1):137–143, ???? 1940. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980574.

Anonymous:1940:PRa

[28] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 103(1): 144–151, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980575.

Anonymous:1940:BMa

[29] Anonymous. Back matter. Journal of the Royal Statistical Society, 103(1):??, ???? 1940. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980576.

Anonymous:1940:FMb

[30] Anonymous. Front matter. Journal of the Royal Statistical Society, 103(2): i–viii, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980413.

Edelberg:1940:FYT

[31] Victor Edelberg. Flexibility of the yield of taxation–some econometric investigations. Journal of the Royal Statistical Society, 103(2):153–190, ???? 1940. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980414.

Allen:1940:USO

[32] R. G. D. Allen. The unemployment situation at the outbreak of War. Journal of the Royal Statistical Society, 103(2): 191–217, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980415.

Rhodes:1940:PMIa

[33] E. C. Rhodes. Population Mathematics– II. Journal of the Royal Statistical Society, 103(2):218–245, ???? 1940. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980416.

Greenwood:1940:SMN

[34] M. Greenwood. A statistical Mare’s nest? Journal of the Royal Statistical Society, 103(2):246–248, ???? 1940. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980417.

Anonymous:1940:CAS

[35] Anonymous. Centenary of the American Statistical Association. Journal of the Royal Statistical Society, 103(2):249, ???? 1940. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980418.

I:1940:BRB

[36] J. O. I. Book review: The Statistical Theory of Estimation, by R. A. Fisher. Journal of the Royal Statistical Society, 103(2):250, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980419.

Cochran:1940:BRB

[37] W. G. Cochran. Book review: Methods of Statistical Analysis, by C. H. Goulden. Journal of the Royal Statistical Society, 103(2):250–251, ???? 1940. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980420.

K:1940:BRBc

[38] M. G. K. Book review: Elements of Statistical Reasoning, by Alan E. Treloar. Journal of the Royal Statistical Society, 103(2):251–252, ???? 1940. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980421.

K:1940:BRBd

[39] M. G. K. Book review: The Variate Difference Method, by Gerhard Tintner. Journal of the Royal Statistical Society, 103(2):252–254, ???? 1940. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980422.

K:1940:BRBe

[40] M. G. K. Book review: A Study of Accident Proneness among Motor Drivers, by E. Farmer; E. G. Chambers. Journal of the Royal Statistical Society, 103(2): 254–256, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980423.

W:1940:BRBd

[41] J. E. W. Book review: Statistical Testing of Business-Cycle Theories. I. A Method and its Application to Investment Activity, by J. Tinbergen; Statistical Testing of Business-Cycle Theories. II. Business Cycles in the United States of America, by J. Tinbergen. Journal of the Royal Statistical Society, 103(2):256–259, ???? 1940. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980424.

G:1940:BRBc

[42] C. O. G. Book review: L’Économie Politique Contemporaine ; La Méthode de l’Économi Politique, by Bertrand Nogaro; L’Intervention de l’État en Matière Économique by Henry Laufenburger; Production Agricole et Économie Rurale by André Garrigou-Lagrange; L’Économie Collectiviste by Robert Mossé; Initiation à l’Analyse Statistique by Léo Dugé de Bernonville. Journal of the Royal Statistical Society, 103(2): 259–261, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980425.

N:1940:BRB

[43] A. M. N. Book review: Profits, Interest and Investment and Other Essays on the Theory of Industrial Fluctuations, by F. A. von Hayek. Journal of the Royal Statistical Society, 103(2):262–263, ???? 1940. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2980426.

G:1940:BRBd

[44] H. W. G. G. Book review: The Economics of Building, by Herbert W. Robinson. Journal of the Royal Statistical Society, 103(2):263–265, ???? 1940. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980427.

M:1940:BRBc

[45] H. W. M. Book review: Prices and Wages in England from the Twelfth to the Nineteenth Century, by William Beveridge. Journal of the Royal Statistical Society, 103(2):265, ???? 1940. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980428.

Anonymous:1940:SNb

[46] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 103(2): 266–274, ???? 1940. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980429.

Anonymous:1940:SEAb

[47] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 103(2): 275–279, ???? 1940. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980430.

Anonymous:1940:LALb

[48] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 103(2):280–284, ???? 1940. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980431.

Anonymous:1940:BMb

[49] Anonymous. Back matter. Journal of the Royal Statistical Society, 103(2):??, ???? 1940. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980432.

Anonymous:1940:FMc

[50] Anonymous. Front matter. Journal of the Royal Statistical Society, 103(3): i–viii, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980451.

Grebenik:1940:SAP

[51] E. Grebenik. Some aspects of population in Bristol. Journal of the Royal Statistical Society, 103(3):285–322, ???? 1940. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980452.

Jones:1940:RRC

[52] J. H. Jones. The report of the Royal Commission on the Distribution of the Industrial Population. Journal of the Royal Statistical Society, 103(3):323– 343, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980453.

Anonymous:1940:WP

[53] Anonymous. Wholesale prices in 1939. Journal of the Royal Statistical Society, 103(3):344–361, ???? 1940. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980454.

Rhodes:1940:PMIb

[54] E. C. Rhodes. Population Mathematics– III. Journal of the Royal Statistical Society, 103(3):362–387, ???? 1940. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980455.

Kendall:1940:MML

[55] M. G. Kendall. On the method of maximum likelihood. Journal of the Royal Statistical Society, 103(3):388–399, ???? 1940. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980456.

H:1940:BRBc

[56] H. O. H. Book review: Statistical Mathematics, by A. C. Aitken. Journal of the Royal Statistical Society, 103(3): 400–401, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980457.

H:1940:BRBd

[57] H. O. H. Book review: Mathematics of Statistics, by J. F. Kenney. Journal of the Royal Statistical Society, 103(3): 401–403, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980458.

B:1940:BRBc

[58] M. S. B. Book review: Statistical Technique in Agricultural Research: A Simple Exposition of Practice and Procedure in Biometry, by D. D. Paterson. Journal of the Royal Statistical Society, 103(3): 403–404, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980459.

B:1940:BRBd

[59] H. S. B. Book review: Graphic Presentation, by Willard C. Brinton. Journal of the Royal Statistical Society, 103(3):404, ???? 1940. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980460.

R:1940:BRBa

[60] E. C. R. Book review: Population Policies and Movements in Europe, by D.

V. Glass. Journal of the Royal Statistical Society, 103(3):405–406, ???? 1940. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980461.

E:1940:BRBb

[61] P. G. E. Book review: The Cameroons and Togoland, by Robert R. Kuczynski. Journal of the Royal Statistical Society, 103(3):406–407, ???? 1940. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980462.

K:1940:BRBf

[62] M. G. K. Book review: Indian Agricultural Statistics. An Introductory Study, by P. J. Thomas; N. Sundararama Sastry. Journal of the Royal Statistical Society, 103(3):407–409, ???? 1940. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980463.

W:1940:BRBe

[63] A. D. W. Book review: Alcohol and the Nation: A Contribution to the Study of the Liquor Problem in the United Kingdom from 1800–1935, by George B. Wilson. Journal of the Royal Statistical Society, 103(3):409–411, ???? 1940. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980464.

B:1940:BRBe

[64] H. S. B. Book review: The Minimum Wage. An International Survey. Journal of the Royal Statistical Society, 103(3): 411–412, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980465.

G:1940:BRBe

[65] C. O. G. Book review: The Incidence of Income Taxes, by Duncan Black. Journal of the Royal Statistical Society, 103 (3):412–413, ???? 1940. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980466.

W:1940:BRBf

[66] J. E. W. Book review: Essays in Monetary Theory, by D. H. Robertson. Journal of the Royal Statistical Society, 103 (3):413–415, ???? 1940. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980467.

W:1940:BRBg

[67] J. E. W. Book review: The Stock Market, Credit and Capital Formation, by Fritz Machlup. Journal of the Royal Statistical Society, 103(3):415–417, ???? 1940. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2980468.

R:1940:BRBb

[68] M. R. Book review: Tin Restriction and Tin Price, by M. J. Schut. Journal of the Royal Statistical Society, 103(3): 417–418, ???? 1940. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980469.

A:1940:BRBc

[69] J. E. A. Book review: The Road and Rail Transport Problem, by H. Osborne Mance; Arthur Salter. Journal of the Royal Statistical Society, 103(3): 418–421, ???? 1940. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980470.

Anonymous:1940:ONPb

[70] Anonymous. Other new publications. Journal of the Royal Statistical Society, 103(3):421–423, ???? 1940. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980471.

Anonymous:1940:SNc

[71] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 103(3): 424–433, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980472.

Anonymous:1940:CNb

[72] Anonymous. Current notes. Journal of the Royal Statistical Society, 103(3):434, ???? 1940. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980473.

Anonymous:1940:SEAc

[73] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 103(3): 435–439, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980474.

Anonymous:1940:LALc

[74] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 103(3):440–443, ???? 1940. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980475.

Anonymous:1940:PRb

[75] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 103(3): 444–449, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980476.

Anonymous:1940:BMc

[76] Anonymous. Back matter. Journal of the Royal Statistical Society, 103(3):??, ???? 1940. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980477.

Anonymous:1940:VI

[77] Anonymous. Volume information. Journal of the Royal Statistical Society, 103 (4):614–vii, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980686.

Anonymous:1940:FMd

[78] Anonymous. Front matter. Journal of the Royal Statistical Society, 103(4): i–viii, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980687.

Macrosty:1940:OTU

[79] Henry W. Macrosty. The overseas trade of the United Kingdom 1930–39. Journal of the Royal Statistical Society, 103(4): 451–490, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980688.

Bowley:1940:SCN

[80] A. L. Bowley. Some constituents of the national income. Journal of the Royal Statistical Society, 103(4):491–518, ???? 1940. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980689.

Anonymous:1940:PM

[81] Anonymous. Proceedings of the meeting. Journal of the Royal Statistical Society, 103(4):519–523, ???? 1940. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980690.

Marley:1940:CSG

[82] Joan G. Marley and H. Campion. Changes in salaries in Great Britain, 1924–1939. Journal of the Royal Statistical Society, 103(4):524–533, ???? 1940. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980691.

Hartley:1940:RAM

[83] H. O. Hartley. Recent advances in mathematical statistics. Journal of the Royal Statistical Society, 103(4):534–560, ???? 1940. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980692.

Anonymous:1940:RCF

[84] Anonymous. Report of the Council for the financial year ending June 18th, 1940, presented at the one hundred and sixth Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts, John Street, Adelphi, W.C. 2, on June 18th, 1940. Journal of the Royal Statistical Society, 103(4):561–571, ???? 1940. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980693.

Anonymous:1940:POH

[85] Anonymous. Proceedings of the one hundred and sixth Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts on Tuesday, June 18th, 1940. Journal of the Royal Statistical Society, 103(4):572, ???? 1940. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980694.

H:1940:BRBe

[86] H. O. H. Book review: Statistical Dictionary of Terms and Symbols, by A. K. Kurtz; H. A. Edgerton. Journal of the Royal Statistical Society, 103(4): 573–574, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980695.

G:1940:BRBf

[87] C. O. G. Book review: The Real Cost of the War, by J. Keith Horsefield. Journal of the Royal Statistical Society, 103(4): 574–575, ???? 1940. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980696.

M:1940:BRBd

[88] H. W. M. Book review: The City: The London Financial Markets, by Barnard Ellinger. Journal of the Royal Statistical Society, 103(4):575–577, ???? 1940. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980697.

M:1940:BRBe

[89] H. W. M. Book review: The Economic History of Steelmaking, 1867–1939: A Study in Competition, by D. L. Burn. Journal of the Royal Statistical Society, 103(4):578–580, ???? 1940. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980698.

M:1940:BRBf

[90] H. W. M. Book review: Growth of American Manufacturing Areas: A Comparative Analysis, with Special Emphasis on Trends in the Pittsburgh District, by Glenn E. McLaughlin. Journal of the Royal Statistical Society, 103(4): 580–581, ???? 1940. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980699.

B:1940:BRBf

[91] G. A. B. Book review: International Monetary Economics, by Michael A. Heilperi. Journal of the Royal Statistical Society, 103(4):581–583, ???? 1940. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980700.

B:1940:BRBg

[92] G. A. B. Book review: Theory of the Just Price, by Rudolf Kaulla. Journal of the Royal Statistical Society, 103(4):583, ???? 1940. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980701.

Anonymous:1940:SNd

[93] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 103(4): 584–592, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980702.

Anonymous:1940:CNc

[94] Anonymous. Current notes. Journal of the Royal Statistical Society, 103(4): 593–595, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980703.

Anonymous:1940:SEAd

[95] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 103(4): 596–598, ???? 1940. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980704.

Anonymous:1940:LALd

[96] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 103(4):599–601, ???? 1940. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980705.

Anonymous:1940:PRL

[97] Anonymous. Periodicals received by the Library. Journal of the Royal Statistical Society, 103(4):602–613, ???? 1940. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980706.

Anonymous:1940:BMd

[98] Anonymous. Back matter. Journal of the Royal Statistical Society, 103(4):??, ???? 1940. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2980707.

Anonymous:1941:VIa

[99] Anonymous. Volume information. Journal of the Royal Statistical Society, 104 (1):v–vii, ???? 1941. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980254.

Anonymous:1941:FM

[100] Anonymous. Front matter. Journal of the Royal Statistical Society, 104(1):i–iii, ???? 1941. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2980255.

Macrosty:1941:ESR

[101] Henry W. Macrosty. Economic statistic– retrospect and prospect. Journal of the Royal Statistical Society, 104(1):1–14, ???? 1941. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2980256.

Rhodes:1941:SCD

[102] E. C. Rhodes. Secular changes in death rates. Journal of the Royal Statistical Society, 104(1):15–42, ???? 1941. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980257.

Kendall:1941:EET

[103] M. G. Kendall. The effect of the elimination of trend on oscillation in timeseries. Journal of the Royal Statistical Society, 104(1):43–52, ???? 1941. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980258.

Anonymous:1941:CLW

[104] Anonymous. Cost of living of the working classes. Journal of the Royal Statistical Society, 104(1):53–58, ???? 1941.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2980259.

I:1941:BRB

[105] J. O. I. Book review: Theory of Probability, by Harold Jeffreys. Journal of the Royal Statistical Society, 104(1):59–64, ???? 1941. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980260.

H:1941:BRBa

[106] H. O. H. Book review: Statistical Calculations for Beginners, by E. G. Chambers. Journal of the Royal Statistical Society, 104(1):64–65, ???? 1941. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980261.

H:1941:BRBb

[107] H. O. H. Book review: Statistical Methods for Medical and Biological Students, by Gunnar Dahlberg. Journal of the Royal Statistical Society, 104(1):65–66, ???? 1941. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980262.

H:1941:BRBc

[108] H. O. H. Book review: A Detailed Proof of the Chi-Square Test of Goodness of Fit, by E. Russell Greenhood. Journal of the Royal Statistical Society, 104 (1):66–67, ???? 1941. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980263.

G:1941:BRBa

[109] E. G. Book review: Education and the Birth Rate, by Grace G. Leybourne; Kenneth White. Journal of the Royal Statistical Society, 104(1):67–68, ???? 1941. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980264.

W:1941:BRB

[110] J. E. W. Book review: Capitalism the Creator, by Carl Snyder. Journal of the Royal Statistical Society, 104(1):68–70, ???? 1941. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2980265.

R:1941:BRBa

[111] H. W. R. Book review: Unemployment and the Unemployed, by H. W. Singer. Journal of the Royal Statistical Society, 104(1):70–72, ???? 1941. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980266.

E:1941:BRBa

[112] P. G. E. Book review: Penal Reform in England: Introductory Essays on some Aspects of English Criminal Policy, by L. Radzinowicz; J. W. Cecil Turner; P. H. Winfield. Journal of the Royal Statistical Society, 104(1):72–73, ???? 1941. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980267.

Anonymous:1941:SNa

[113] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 104 (1):74–82, ???? 1941. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980268.

Anonymous:1941:CNa

[114] Anonymous. Current notes. Journal of the Royal Statistical Society, 104(1): 83–84, ???? 1941. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980269.

Anonymous:1941:HWM

[115] Anonymous. H. W. M. Journal of the Royal Statistical Society, 104(1):85–90,

???? 1941. URL http://www.jstor.
org/stable/2980270.

Anonymous:1941:JB

[116] Anonymous. James Bonar. Journal of the Royal Statistical Society, 104(1): 91–94, ???? 1941. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980271.

Anonymous:1941:RP

[117] Anonymous. Raymond Pearl. Journal of the Royal Statistical Society, 104(1): 94–96, ???? 1941. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980272.

Anonymous:1941:SEAa

[118] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 104(1): 97–99, ???? 1941. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980273.

Anonymous:1941:LALa

[119] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 104(1):100–102, ???? 1941. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980274.

Anonymous:1941:PRa

[120] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 104(1): 103–110, ???? 1941. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980275.

Kendall:1941:FBA

[121] M. G. Kendall. The financing of British agriculture. Journal of the Royal Statistical Society, 104(2):111–142, ???? 1941.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2980031.

Anonymous:1941:PFO

[122] Anonymous. Proceedings of the first ordinary meeting of the session 1940– 41 held on April 15th, 1941, the President, Mr. H. Leak, in the Chair. Journal of the Royal Statistical Society, 104(2): 143–145, ???? 1941. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980032.

Greenwood:1941:DV

[123] M. Greenwood, W. J. Martin, and W. T. Russell. Deaths by violence, 1837–1937. Journal of the Royal Statistical Society, 104(2):146–171, ???? 1941. ISSN 0952-

8385. URL http://www.jstor.org/

stable/2980033. See erratum [138].

Turin:1941:SOP

[124] S. P. Turin. Some observations on the population of Soviet Russia at the census of January 17th, 1939. Journal of the Royal Statistical Society, 104(2): 172–174, ???? 1941. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980034.

K:1941:BRBa

[125] M. G. K. Book review: Statistical Procedures and their Mathematical Bases, by C. C. Peters; W. R. Van Voorhis. Journal of the Royal Statistical Society, 104(2):175–176, ???? 1941. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980035.

K:1941:BRBb

[126] M. G. K. Book review: The Development of Mathematics, by E. T. Bell. Journal of the Royal Statistical Society, 104(2):176–177, ???? 1941. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980036.

K:1941:BRBc

[127] M. G. K. Book review: Business Cycles: A Theoretical, Historical and Statistical Analysis of the Capitalist Process, by Joseph A. Schumpeter. Journal of the Royal Statistical Society, 104(2): 177–180, ???? 1941. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980037.

R:1941:BRBb

[128] M. R. Book review: Studies in War Economics, by The International Labour Office. Journal of the Royal Statistical Society, 104(2):180–181, ???? 1941. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980038.

G:1941:BRBb

[129] C. O. G. Book review: Studies in Economic Warfare, by D. T. Jack. Journal of the Royal Statistical Society, 104(2): 181–182, ???? 1941. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980039.

R:1941:BRBc

[130] E. C. R. Book review: Juvenile Delinquency, a Comparative Study of the Position in Liverpool and England and Wales, by J. H. Bogot. Journal of the Royal Statistical Society, 104(2):182– 183, ???? 1941. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980040.

C:1941:BRB

[131] L. R. C. Book review: The Excess Profits Tax: A Practical Exposition, by H. E. Seed. Journal of the Royal Statistical Society, 104(2):183–184, ???? 1941. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980041.

Anonymous:1941:SNb

[132] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 104(2): 185–190, ???? 1941. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980042.

Anonymous:1941:CNb

[133] Anonymous. Current notes. Journal of the Royal Statistical Society, 104(2): 191–192, ???? 1941. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980043.

Bowley:1941:LS

[134] A. L. Bowley. Lord Stamp. Journal of the Royal Statistical Society, 104(2): 193–196, ???? 1941. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980044.

Anonymous:1941:SEAb

[135] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 104(2): 197–201, ???? 1941. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980045.

Anonymous:1941:LALb

[136] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 104(2):202–203, ???? 1941. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980046.

Anonymous:1941:PRb

[137] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 104(2): 204–208, ???? 1941. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980047.

Anonymous:1941:EDV

[138] Anonymous. Erratum: Deaths by violence, 1837–1937. Journal of the Royal Statistical Society, 104(3):??, ???? 1941. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980048. See [123].

Wilson:1941:FIT

[139] Duncan Wilson. Factory inspection:

A thirty-five years retrospect. Journal of the Royal Statistical Society, 104(3): 209–234, ???? 1941. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980049.

George:1941:BPF

[140] C. Oswald George. British public finance in peace and war. Journal of the Royal Statistical Society, 104(3):235–280, ???? 1941. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980050.

Kendall:1941:RBC

[141] M. G. Kendall. The relationship between correlation formulae and elliptic functions. Journal of the Royal Statistical Society, 104(3):281–283, ???? 1941. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980051.

H:1941:BRBd

[142] H. O. H. Book review: Field Trials: Their Lay-Out and Statistical Analysis, by John Wishart. Journal of the Royal Statistical Society, 104(3):284, ???? 1941. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980052.

H:1941:BRBe

[143] H. O. H. Book review: Tests of Significance, What They Mean and How to Use Them, by John H. SMith. Journal of the Royal Statistical Society, 104(3):285, ???? 1941. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980053.

E:1941:BRBb

[144] P. G. E. Book review: Studies in American Demography, by Walter F. Willcox. Journal of the Royal Statistical Society, 104(3):285–286, ???? 1941. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980054.

G:1941:BRBc

[145] E. G. Book review: Social and Economic Aspects of Swedish Population Movements 1750–1933, by Dorothy Swaine Thomas. Journal of the Royal Statistical Society, 104(3):286–287, ???? 1941. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980055.

R:1941:BRBd

[146] H. W. R. Book review: Migration to and from the British Isles: Problems and Policies, by R. S. Walshaw. Journal of the Royal Statistical Society, 104(3): 287–288, ???? 1941. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980056.

G:1941:BRBd

[147] C. O. G. Book review: Public Investment and the Trade Cycle in Great Britain, by R. F. Bretherton; F. A. Burchardt; R. S. G. Rutherford. Journal of the Royal Statistical Society, 104(3): 288–290, ???? 1941. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980057.

R:1941:BRBe

[148] H. W. R. Book review: The Labour Cost of the World War to Great Britain, 1914–22: A Statistical Analysis, by N. B. Dearle. Journal of the Royal Statistical Society, 104(3):290–292, ???? 1941. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980058.

R:1941:BRBf

[149] E. C. R. Book review: Social Aspects of Crime in England between the Wars, by Hermann Mannheim. Journal of the Royal Statistical Society, 104(3): 292–293, ???? 1941. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980059.

Anonymous:1941:SNc

[150] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 104(3): 294–301, ???? 1941. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980060.

Anonymous:1941:CNc

[151] Anonymous. Current notes. Journal of the Royal Statistical Society, 104(3): 302–303, ???? 1941. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980061.

Anonymous:1941:SEAc

[152] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 104(3): 304–307, ???? 1941. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980062.

Anonymous:1941:LALc

[153] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 104(3):308–310, ???? 1941. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980063.

Anonymous:1941:VIb

[154] Anonymous. Volume information. Journal of the Royal Statistical Society, 104 (4):411–413, ???? 1941. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980119.

Stocks:1941:DSF

[155] Percy Stocks. Diphteria and scarlet fever incidence during the dispersal of 1939– 40. Journal of the Royal Statistical Society, 104(4):311–345, ???? 1941. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980120.

Anonymous:1941:WP

[156] Anonymous. Wholesale prices in 1940. Journal of the Royal Statistical Society, 104(4):346–362, ???? 1941. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980121.

Anonymous:1941:RCF

[157] Anonymous. Report of the Council for the financial year ended December 31st, 1940, and for the sessional year ending June 24th, 1941, presented at the one hundred and seventh Annual General Meeting of the Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine, W.C. 1, on June 24th, 1941. Journal of the Royal Statistical Society, 104(4):363–371, ???? 1941. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2980122.

Anonymous:1941:POH

[158] Anonymous. Proceedings of the one hundred and seventh Annual General Meeting of the Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine on Tuesday, June 24th, 1941. Journal of the Royal Statistical Society, 104(4):372, ???? 1941. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980123.

K:1941:BRBd

[159] M. G. K. Book review: The Second Yearbook of Research and Statistical Methodology Books and Reviews, by Oskar Krisen Buros. Journal of the Royal Statistical Society, 104(4):373–376, ???? 1941. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980124.

S:1941:BRB

[160] R. S. Book review: The Theory and Measurement of Demand, by Henry Schultz. Journal of the Royal Statistical Society, 104(4):376–378, ???? 1941. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980125.

G:1941:BRBe

[161] C. O. G. Book review: The Taxation of War Wealth, by J. R. Hicks; U. K. Hicks; L. Rostas. Journal of the Royal Statistical Society, 104(4):379–380, ???? 1941. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980126.

R:1941:BRBg

[162] T. M. R. Book review: Price Research in the Steel and Petroleum Industries. Journal of the Royal Statistical Society, 104(4):381–382, ???? 1941. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980127.

J:1941:BRB

[163] R. H. J. Book review: Methods of Assessing the Physical Fitness of Children, by Rachel M. Jenss; Susan P. Souther.

Journal of the Royal Statistical Society, 104(4):382–384, ???? 1941. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980128.

Anonymous:1941:SNd

[164] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 104(4): 385–393, ???? 1941. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980129.

Anonymous:1941:CNd

[165] Anonymous. Current notes. Journal of the Royal Statistical Society, 104(4): 393–395, ???? 1941. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980130.

Anonymous:1941:SEAd

[166] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 104(4): 396–399, ???? 1941. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980131.

Anonymous:1941:LALd

[167] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 104(4):400–402, ???? 1941. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980132.

Anonymous:1941:PRL

[168] Anonymous. Periodicals received by the Library. Journal of the Royal Statistical Society, 104(4):403–410, ???? 1941. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980133.

Anonymous:1942:FMa

[169] Anonymous. Front matter. Journal of the Royal Statistical Society, 105(1):i–vi, ???? 1942. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980192.

Greenwood:1942:BLL

[170] Major Greenwood. British loss of life in the Wars of 1794–1815 and in 1914– 1918. Journal of the Royal Statistical Society, 105(1):1–16, ???? 1942. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980193.

Beveridge:1942:ERA

[171] William Beveridge and Mr. R. G. Glenday. Economic reconstruction after the war: A discussion before the Royal Statistical Society, on March 17th, 1942, the President, Sir William Beveridge, K.C.B., in the chair. Journal of the Royal Statistical Society, 105(1):17–35, ???? 1942. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980194.

Thompson:1942:SSR

[172] R. J. Thompson. Some suggestions in regard to our agricultural statistics. Journal of the Royal Statistical Society, 105 (1):36–41, ???? 1942. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980195.

Moyal:1942:APD

[173] J. E. Moyal. Approximate probability distribution functions for the sum of two independent variates. Journal of the Royal Statistical Society, 105(1):42–43, ???? 1942. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980196.

K:1942:BRBa

[174] M. G. K. Book review: A Significance Test for Time Series, by W. Allen Wallis; Geoffrey H. Moore. Journal of the Royal Statistical Society, 105(1):44–45, ???? 1942. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2980197.

G:1942:BRBa

[175] R. F. G. Book review: Poverty and Progress: A Second Social Survey of York, by R. Seebohm Rowntree. Journal of the Royal Statistical Society, 105 (1):45–48, ???? 1942. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980198.

R:1942:BRBa

[176] F. E. R. Book review: Food Production in Western Europe, by P. Lamartine Yates. Journal of the Royal Statistical Society, 105(1):48–49, ???? 1942. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980199.

R:1942:BRBb

[177] T. M. R. Book review: Railways in Modern India, by R. D. Tiwari; Railway Rates Policy. by R. D. Tiwari. Journal of the Royal Statistical Society, 105 (1):49–50, ???? 1942. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980200.

Anonymous:1942:SNa

[178] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 105 (1):51–55, ???? 1942. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980201.

Anonymous:1942:CNa

[179] Anonymous. Current notes. Journal of the Royal Statistical Society, 105(1):55, ???? 1942. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2980202.

B:1942:JCS

[180] P. E. B. J. Calvert Spensley, O.B.E.

Journal of the Royal Statistical Society, 105(1):56–57, ???? 1942. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980203.

Anonymous:1942:SEAa

[181] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 105(1): 58–60, ???? 1942. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980204.

Anonymous:1942:LALa

[182] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 105(1):61–64, ???? 1942. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980205.

Anonymous:1942:PRa

[183] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 105 (1):65–67, ???? 1942. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980206.

Anonymous:1942:BMa

[184] Anonymous. Back matter. Journal of the Royal Statistical Society, 105(1):??, ???? 1942. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980207.

Anonymous:1942:FMb

[185] Anonymous. Front matter. Journal of the Royal Statistical Society, 105(2):i–vi, ???? 1942. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980608.

Kendall:1942:FS

[186] M. G. Kendall. On the future of statistics. Journal of the Royal Statistical Society, 105(2):69–91, ???? 1942. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980609.

Schwartz:1942:SFS

[187] G. L. Schwartz. The significance of fiscal statistics. Journal of the Royal Statistical Society, 105(2):92–114, ???? 1942. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980610.

Irwin:1942:DWE

[188] J. O. Irwin. On the distribution of a weighted estimate of variance and on analysis of variance in certain cases of unequal weighting. Journal of the Royal Statistical Society, 105(2):115–118, ???? 1942. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980611.

Kendall:1942:NER

[189] M. G. Kendall. Note on the estimation of a ranking. Journal of the Royal Statistical Society, 105(2):119–121, ???? 1942. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980612.

Collet:1942:PPW

[190] Clara E. Collet. The present position of women in industry. Journal of the Royal Statistical Society, 105(2):122–124, ???? 1942. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980613.

K:1942:BRBb

[191] M. G. K. Book review: The Analysis of Economic Time Series, by H. T. Davis. Journal of the Royal Statistical Society, 105(2):125–127, ???? 1942. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980614.

A:1942:BRB

[192] F. J. A. Book review: Guide for Quality Control and Control Chart Method of Analysing Data. ; Quality Control Charts, by B. P. Dudding; W. J. Jennett. Journal of the Royal Statistical Society, 105(2):127–129, ???? 1942. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980615.

W:1942:BRBa

[193] J. E. W. Book review: National Income and its Composition, 1919–1938, by Simon Kuznets; Lillian Epstein; Elizabeth Jenks. Journal of the Royal Statistical Society, 105(2):129–132, ???? 1942. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980616.

G:1942:BRBb

[194] C. O. G. Book review: Standards of Living, Wages and Prices, by H. Belshaw. Journal of the Royal Statistical Society, 105(2):132–133, ???? 1942. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980617.

R:1942:BRBc

[195] T. M. R. Book review: The Theory of Incidence of Sales Taxation, by John F. Due. Journal of the Royal Statistical Society, 105(2):133–134, ???? 1942. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980618.

B:1942:BRBa

[196] G. A. B. Book review: The Economic Effects of Advertising, by Neil H. Borden. Journal of the Royal Statistical Society, 105(2):134–135, ???? 1942. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980619.

Anonymous:1942:SNb

[197] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 105(2): 136–141, ???? 1942. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980620.

Anonymous:1942:CNb

[198] Anonymous. Current notes. Journal of the Royal Statistical Society, 105(2): 141–143, ???? 1942. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980621.

Anonymous:1942:SAW

[199] Anonymous. Sir Alfred William Flux, C.B. Journal of the Royal Statistical Society, 105(2):144–147, ???? 1942. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980622.

Anonymous:1942:SEAb

[200] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 105(2): 148–149, ???? 1942. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980623.

Anonymous:1942:LALb

[201] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 105(2):150–152, ???? 1942. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980624.

Anonymous:1942:PRb

[202] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 105(2): 153–157, ???? 1942. ISSN 0952-8385.

URL http://www.jstor.org/stable/

2980625. See corrigenda [205].

Anonymous:1942:BMb

[203] Anonymous. Back matter. Journal of the Royal Statistical Society, 105(2):??, ???? 1942. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980626.

Anonymous:1942:FMc

[204] Anonymous. Front matter. Journal of the Royal Statistical Society, 105(3):??, ???? 1942. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979984.

Anonymous:1942:CPR

[205] Anonymous. Corrigenda: Periodical returns. Journal of the Royal Statistical Society, 105(3):??, ???? 1942. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979985. See [202].

Massey:1942:EBM

[206] Philip Massey. The expenditure of 1,360 British middle-class households in 1938– 39. Journal of the Royal Statistical Society, 105(3):159–196, ???? 1942. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979986.

Daniel:1942:SEC

[207] G. H. Daniel. Social and economic conditions and the incidence of rheumatic heart disease. Journal of the Royal Statistical Society, 105(3):197–212, ???? 1942. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979987.

Geary:1942:EMP

[208] R. C. Geary. The estimation of many parameters. Journal of the Royal Statistical Society, 105(3):213–217, ???? 1942.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2979988. See corri-

genda [243].

T:1942:LUR

[209] R. J. T. Land utilization in rural areas. Journal of the Royal Statistical Society, 105(3):218–219, ???? 1942. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979989.

Anonymous:1942:WP

[210] Anonymous. Wholesale prices in 1941. Journal of the Royal Statistical Society, 105(3):220–235, ???? 1942. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979990.

W:1942:BRBb

[211] J. E. W. Book review: The International Gold Standard Re-Interpreted. 1914–1934, by William Adams Brown, jun. Journal of the Royal Statistical Society, 105(3):236–238, ???? 1942. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979991.

N:1942:BRB

[212] J. L. N. Book review: Fluctuations in Income and Employment, by T. Wilson. Journal of the Royal Statistical Society, 105(3):238–240, ???? 1942. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979992.

R:1942:BRBd

[213] H. W. R. Book review: Three Aspects of Labour Dynamics, by W. S. Woytinsky. Journal of the Royal Statistical Society, 105(3):240–241, ???? 1942. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979993.

R:1942:BRBe

[214] T. M. R. Book review: Consumer Instalment Credit and Economic Fluctuations, by Gottfried Haberler. Journal of the Royal Statistical Society, 105(3): 241–242, ???? 1942. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979994.

R:1942:BRBf

[215] T. M. R. Book review: Retail Trade Associations, by Hermann Levy. Journal of the Royal Statistical Society, 105(3): 242–243, ???? 1942. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979995.

R:1942:BRBg

[216] E. D. R. Book review: Measurement of Trees, by Reginald Davey. Journal of the Royal Statistical Society, 105(3): 243–244, ???? 1942. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979996.

G:1942:BRBc

[217] C. O. G. Book review: An Essay on Marxian Economics, by Joan Robinson. Journal of the Royal Statistical Society, 105(3):244–245, ???? 1942. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979997.

Anonymous:1942:SNc

[218] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 105(3): 246–251, ???? 1942. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979998.

Anonymous:1942:CNc

[219] Anonymous. Current notes. Journal of the Royal Statistical Society, 105(3):

251–253, ???? 1942. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979999.

Anonymous:1942:SEAc

[220] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 105(3): 254–255, ???? 1942. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980000.

Anonymous:1942:LALc

[221] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 105(3):256–258, ???? 1942. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980001.

Anonymous:1942:BMc

[222] Anonymous. Back matter. Journal of the Royal Statistical Society, 105(3):??, ???? 1942. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2980002.

Anonymous:1942:VI

[223] Anonymous. Volume information. Journal of the Royal Statistical Society, 105 (4):376–iv, ???? 1942. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980433.

Anonymous:1942:FMd

[224] Anonymous. Front matter. Journal of the Royal Statistical Society, 105(4):??, ???? 1942. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2980434.

Stocks:1942:MWC

[225] Percy Stocks. Measles and whoopingcough incidence before and during the dispersal of 1939–41. Journal of the Royal Statistical Society, 105(4):259– 291, ???? 1942. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980435.

Neyman:1942:BIS

[226] J. Neyman. Basic ideas and some recent results of the theory of testing statistical hypotheses. Journal of the Royal Statistical Society, 105(4):292–327, ???? 1942. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980436.

Darke:1942:SGP

[227] W. F. Darke. A short guide to prewar English outdoor vegetable statistics. Journal of the Royal Statistical Society, 105(4):328–335, ???? 1942. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980437.

Gibbon:1942:BR

[228] Gwilym Gibbon. The Beveridge report.

Journal of the Royal Statistical Society, 105(4):336–340, ???? 1942. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980438.

Anonymous:1942:RCF

[229] Anonymous. Report of the Council for the financial year ended December 31st, 1941, and for the sessional year ending June 16th, 1942, presented at the one hundred and eighth Annual General Meeting of the Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine, W.C. 1, on June 16th, 1942. Journal of the Royal Statistical Society, 105(4):341–348, ???? 1942. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980439.

Anonymous:1942:POH

[230] Anonymous. Proceedings of the one hundred and eighth general meeting of the Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine on Tuesday, June 16th, 1942. Journal of the Royal Statistical Society, 105(4):349, ???? 1942. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980440.

E:1942:BRB

[231] W. P. E. Book review: The Fundamental Principles of Mathematical Statistics, by Hugh H. Wolfenden. Journal of the Royal Statistical Society, 105(4): 350–351, ???? 1942. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980441.

G:1942:BRBd

[232] C. O. G. Book review: The Burden of British Taxation, by G. Findlay Shirras; L. Rostas. Journal of the Royal Statistical Society, 105(4):351–352, ???? 1942. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980442.

S:1942:BRB

[233] E. C. S. Book review: The Economics of 1960, by Colin Clark. Journal of the Royal Statistical Society, 105(4):352– 353, ???? 1942. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980443.

B:1942:BRBb

[234] G. A. B. Book review: The Theory of Money, by F. Henry. Journal of the Royal Statistical Society, 105(4): 353–354, ???? 1942. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980444.

G:1942:BRBe

[235] H. G. Book review: Review of Housing Before the War and After, by M. J. Elsas. Journal of the Royal Statistical Society, 105(4):354–355, ???? 1942. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980445.

Anonymous:1942:SNd

[236] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 105(4): 356–361, ???? 1942. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980446.

Anonymous:1942:SEAd

[237] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 105(4): 362–364, ???? 1942. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980447.

Anonymous:1942:LALd

[238] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 105(4):365–369, ???? 1942. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980448.

Anonymous:1942:PRL

[239] Anonymous. Periodicals received by the Library. Journal of the Royal Statistical Society, 105(4):370–375, ???? 1942. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980449.

Anonymous:1942:BMd

[240] Anonymous. Back matter. Journal of the Royal Statistical Society, 105(4):??, ???? 1942. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980450.

Anonymous:1943:VIa

[241] Anonymous. Volume information. Journal of the Royal Statistical Society, 106 (1):iii–iv, ???? 1943. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980236.

Anonymous:1943:FM

[242] Anonymous. Front matter. Journal of the Royal Statistical Society, 106(1):i–ii, ???? 1943. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980237.

Geary:1943:CEM

[243] R. C. Geary. Corrigenda: Estimation of many parameters. Journal of the Royal Statistical Society, 106(1):??, ???? 1943. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980238. See

[208].

Elderton:1943:MAM

[244] William P. Elderton and M. E. Ogborn. The mortality of adult males since the middle of the Eighteenth Century as shown by the experience of life assurance companies. Journal of the Royal Statistical Society, 106(1):1–31, ???? 1943. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980239.

Thompson:1943:FA

[245] R. J. Thompson. The future of agriculture. Journal of the Royal Statistical Society, 106(1):32–50, ???? 1943. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980240.

Lomax:1943:RBE

[246] K. S. Lomax. The relationship between expenditure per head and size of population of county boroughs in England and Wales. Journal of the Royal Statistical Society, 106(1):51–59, ???? 1943. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980241.

Leak:1943:VIP

[247] H. Leak. Volume of industrial production. Journal of the Royal Statistical Society, 106(1):60–63, ???? 1943. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980242. See [350].

Dudding:1943:IAG

[248] Bernard P. Dudding. The industrial applications group of the Royal Statistical Society: First session, 1942–43. Journal of the Royal Statistical Society, 106 (1):64–67, ???? 1943. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980243.

Finney:1943:BRB

[249] D. J. Finney. Book review: Statistical Analysis in Biology, by K. Mather. Journal of the Royal Statistical Society, 106 (1):68–69, ???? 1943. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980244.

E:1943:BRBa

[250] M. J. E. Book review: The Population of Bristol, by H. A. Shannon; E. Grebenik. Journal of the Royal Statistical Society, 106(1):69–70, ???? 1943. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980245.

B:1943:BRBa

[251] G. A. B. Book review: The Social Framework: An Introduction to Economics, by J. R. Hicks. Journal of the Royal Statistical Society, 106(1):70–71, ???? 1943. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980246.

C:1943:BRB

[252] C. E. C. Book review: A History of the Straw Hat Industry, by John G. Dony.

Journal of the Royal Statistical Society, 106(1):71, ???? 1943. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980247.

H:1943:BRB

[253] W. M. H. Book review: Inter-American Statistical Yearbook, 1942, by Raul C. Migone. Journal of the Royal Statistical Society, 106(1):71–72, ???? 1943. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980248.

Anonymous:1943:SNa

[254] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 106 (1):73–77, ???? 1943. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980249.

G:1943:SAN

[255] M. G. Sir Arthur Newsholme, K.C.B. Journal of the Royal Statistical Society, 106(1):78, ???? 1943. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980250.

Anonymous:1943:SEAa

[256] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 106(1): 79–81, ???? 1943. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980251.

Anonymous:1943:LALa

[257] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 106(1):82–87, ???? 1943. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980252.

Anonymous:1943:PRa

[258] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 106 (1):88–90, ???? 1943. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980253.

Kendall:1943:OME

[259] M. G. Kendall. Oscillatory movements in English agriculture. Journal of the Royal Statistical Society, 106(2):91–124, ???? 1943. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2980373.

Herdan:1943:LAR

[260] G. Herdan. The logical and analytical relationship between the theory of accidents and factor analysis. Journal of the Royal Statistical Society, 106(2): 125–142, ???? 1943. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980374.

Irwin:1943:TVW

[261] J. O. Irwin. A table of the variance of \sqrt{x} when x has a Poisson distribution. Journal of the Royal Statistical Society, 106(2):143–144, ???? 1943. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980375.

Anonymous:1943:MOS

[262] Anonymous. Memorandum on official statistics. Journal of the Royal Statistical Society, 106(2):145–166, ???? 1943. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980376.

Bowley:1943:VIP

[263] A. L. Bowley. The volume of industrial production. Journal of the Royal Statistical Society, 106(2):167–170, ????

1943. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980377.

K:1943:BRB

[264] M. G. K. Book review: Poisson’s Exponential Binomial Limit, by E. C. Molina. Journal of the Royal Statistical Society, 106(2):171, ???? 1943. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980378.

G:1943:BRBa

[265] E. G. Book review: Birth, Poverty and Wealth. A Study of Infant Mortality, by Richard M. Titmuss. Journal of the Royal Statistical Society, 106(2): 171–172, ???? 1943. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980379.

G:1943:BRBb

[266] R. F. G. Book review: Young Offenders: An Enquiry into Juvenile Delinquency, by A. M. Carr-Saunders; Hermann Mannheim; E. C. Rhodes. Journal of the Royal Statistical Society, 106(2): 172–173, ???? 1943. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980380.

R:1943:BRBa

[267] F. E. R. Book review: The Rice Economy of Monsoon Asia, by V. D. Wickizer; M. K. Bennett. Journal of the Royal Statistical Society, 106(2):173– 174, ???? 1943. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980381.

B:1943:BRBb

[268] D. S. B. Book review: Standardized Accounting in Germany, by H. W. Singer. Journal of the Royal Statistical Society, 106(2):174–176, ???? 1943. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980382.

Anonymous:1943:SNb

[269] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 106(2): 177–183, ???? 1943. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980383.

Anonymous:1943:CNa

[270] Anonymous. Current notes. Journal of the Royal Statistical Society, 106(2): 184–186, ???? 1943. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980384.

Anonymous:1943:SEAb

[271] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 106(2): 187–190, ???? 1943. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980385.

Anonymous:1943:LALb

[272] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 106(2):191–195, ???? 1943. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980386.

Anonymous:1943:PRb

[273] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 106(2): 196–200, ???? 1943. ISSN 0952-8385.

URL http://www.jstor.org/stable/

2980387. See corrigenda [274].

Anonymous:1943:CPR

[274] Anonymous. Corrigenda: Periodical returns. Journal of the Royal Statistical Society, 106(3):??, ???? 1943. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979963. See [273].

Crump:1943:PWI

[275] Norman Crump. Post-war international monetary plans: A discussion before the Royal Statistical Society, on May 18th, 1943, Mr. H. Leak, C.B.E., in the Chair. Journal of the Royal Statistical Society, 106(3):201–213, ???? 1943. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979964.

Shirras:1943:MEB

[276] G. Findlay Shirras. Methods of estimating the burden of taxation. Journal of the Royal Statistical Society, 106(3): 214–249, ???? 1943. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979965.

Anonymous:1943:WP

[277] Anonymous. Wholesale prices in 1942. Journal of the Royal Statistical Society, 106(3):250–265, ???? 1943. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979966.

Simpson:1943:TCS

[278] Harold Simpson. On a theorem concerning sampling. Journal of the Royal Statistical Society, 106(3):266–267, ???? 1943. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979967. See

comments [315].

Woodcock:1943:PPS

[279] Fred Woodcock. The price of provisions and some social consequences in Worcestershire in the Eighteenth and Nineteenth Centuries. Journal of the Royal Statistical Society, 106(3):268–272, ????

1943. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979968. See

corrigendum [296].

Thompson:1943:UNC

[280] R. J. Thompson. The United Nations Conference on Food and Agriculture. Journal of the Royal Statistical Society, 106(3):273–276, ???? 1943. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979969.

Anonymous:1943:LRS

[281] Anonymous. The library of the Royal Statistical Society. Journal of the Royal Statistical Society, 106(3):277, ???? 1943. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979970.

I:1943:BRB

[282] L. I. Book review: “Student’s” Collected Papers. Journal of the Royal Statistical Society, 106(3):278–279, ???? 1943. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979971.

B:1943:BRBc

[283] M. S. B. Book review: Mathematical Statistics, by S. S. Wilks. Journal of the Royal Statistical Society, 106(3): 279–280, ???? 1943. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979972.

R:1943:BRBb

[284] F. E. R. Book review: The American-Born in Canada. A Statistical Interpretation, by R. H. Coats; M. C. Maclean. Journal of the Royal Statistical Society, 106(3):280–281, ???? 1943. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979973.

W:1943:BRB

[285] J. E. W. Book review: The Theory of Prices. A Re-Examination of the Central Problems of Monetary Theory, by Arthur W. Marget. Journal of the Royal Statistical Society, 106(3):281–283, ???? 1943. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979974.

L:1943:BRB

[286] K. S. L. Book review: Standards of Local Expenditure, by J. R. Hicks; U. K. Hicks. Journal of the Royal Statistical Society, 106(3):284–285, ???? 1943. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979975.

R:1943:BRBc

[287] F. E. R. Book review: Modern Commercial Policy, by R. D. Tiwari. Journal of the Royal Statistical Society, 106(3): 285–286, ???? 1943. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979976.

B:1943:BRBd

[288] G. A. B. Book review: Accounts Receivable Financing, by Raymond J. Saulnier; Neil H. Jacoby. Journal of the Royal Statistical Society, 106(3):287, ???? 1943. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2979977.

Anonymous:1943:SNc

[289] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 106(3): 288–291, ???? 1943. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979978.

Anonymous:1943:CNb

[290] Anonymous. Current notes. Journal of the Royal Statistical Society, 106(3):

291–292, ???? 1943. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979979.

R:1943:PJH

[291] E. C. R. Professor John Hilton. Journal of the Royal Statistical Society, 106(3): 293–294, ???? 1943. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979980.

S:1943:LM

[292] G. F. S. Lord Meston 1865–1943. Journal of the Royal Statistical Society, 106 (3):294–296, ???? 1943. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979981.

Anonymous:1943:SEAc

[293] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 106(3): 297–300, ???? 1943. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979982.

Anonymous:1943:LALc

[294] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 106(3):301–303, ???? 1943. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979983.

Anonymous:1943:VIb

[295] Anonymous. Volume information. Journal of the Royal Statistical Society, 106 (4):399–402, ???? 1943. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980478.

Anonymous:1943:CPP

[296] Anonymous. Corrigenda: The price of provisions and some social consequences in Worcestershire in the Eighteenth and Nineteenth Centuries. Journal of the Royal Statistical Society, 106(4):??, ???? 1943. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980479. See

[279].

Beveridge:1943:SSS

[297] William Beveridge. Social security: Some trans-Atlantic comparisons. Journal of the Royal Statistical Society, 106 (4):305–332, ???? 1943. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980480.

Vajda:1943:AAC

[298] S. Vajda. The algebraic analysis of contingency tables. Journal of the Royal Statistical Society, 106(4):333–342, ???? 1943. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980481.

Sandon:1943:CCS

[299] Frank Sandon. Control charts in script assessment in large written examinations. Journal of the Royal Statistical Society, 106(4):343–348, ???? 1943. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980482.

Baykov:1943:NTP

[300] Alexander Baykov. A note on the trend of population and the labour problems of the U.S.S.R. Journal of the Royal Statistical Society, 106(4):349–359, ???? 1943. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980483.

Dyson:1943:NK

[301] F. J. Dyson. A note on kurtosis. Journal of the Royal Statistical Society, 106(4): 360–361, ???? 1943. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980484.

Anonymous:1943:RCF

[302] Anonymous. Report of the Council for the financial year ended December 31st, 1942, and for the sessional year ending June 16th, 1943, presented at the one hundred and ninth Annual General Meeting of the Royal Statistical Society, held at the house of the Royal Society of Arts, Adelphi, W.C. 2, on June 22nd, 1943. Journal of the Royal Statistical Society, 106(4):362–371, ???? 1943. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980485.

Anonymous:1943:POH

[303] Anonymous. Proceedings of the one hundred and ninth Annual General Meeting of the Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine on Tuesday, June 22nd, 1943. Journal of the Royal Statistical Society, 106(4):372, ???? 1943. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980486.

G:1943:BRBc

[304] M. G. Book review: Statistics, by L. H.

C. Tippett. Journal of the Royal Statistical Society, 106(4):373, ???? 1943. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980487.

A:1943:BRB

[305] J. E. A. Book review: The Future Population of Europe and the Soviet Union; Population Projections 1940– 70, by Frank W. Notestein. Journal of the Royal Statistical Society, 106(4): 373–375, ???? 1943. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2980488.

G:1943:BRBd

[306] E. G. Book review: The Displacement of Population in Europe, by Eugene M. Kulischer. Journal of the Royal Statistical Society, 106(4):375–376, ???? 1943. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2980489.

E:1943:BRBb

[307] M. J. E. Book review: The Journey to Work, its Significance for Industrial and Community Life, by Kate K. Liepmann. Journal of the Royal Statistical Society, 106(4):376–377, ???? 1943. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980490.

Anonymous:1943:SNd

[308] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 106(4): 378–384, ???? 1943. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980491.

Anonymous:1943:SEAd

[309] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 106(4): 385–388, ???? 1943. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2980492.

Anonymous:1943:LALd

[310] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 106(4):389–392, ???? 1943. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980493.

Anonymous:1943:PRL

[311] Anonymous. Periodicals received by the Library. Journal of the Royal Statistical Society, 106(4):393–398, ???? 1943.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2980494.

Anonymous:1944:FMa

[312] Anonymous. Front matter. Journal of the Royal Statistical Society, 107(1):??, ???? 1944. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2981351.

Snow:1944:ICI

[313] E. C. Snow. The international comparison of industrial output. Journal of the Royal Statistical Society, 107(1):1–55, ???? 1944. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2981352.

Geary:1944:ETH

[314] R. C. Geary. Extension of a theorem by Harald Cramér on the frequency distribution of the quotient of two variables. Journal of the Royal Statistical Society, 107(1):56–57, ???? 1944. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981353.

Kempthorne:1944:CNT

[315] O. Kempthorne. Comments on the note “On a Theorem Concerning Sampling”. Journal of the Royal Statistical Society, 107(1):58, ???? 1944. ISSN 0952-8385.

URL http://www.jstor.org/stable/

2981354. See [278].

Morrell:1944:NWH

[316] A. J. H. Morrell. Note on Wilson and Hilferty’s approximation to the χ^{2}- distribution. Journal of the Royal Statistical Society, 107(1):59, ???? 1944. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981355.

Dudding:1944:IAG

[317] Bernard P. Dudding. The industrial applications group of the Royal Statistical Society: Second session, 1943–44. Journal of the Royal Statistical Society, 107 (1):60–63, ???? 1944. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981356.

G:1944:BRBa

[318] C. O. G. Book review: American Agriculture, 1899–1939, by Harold Barger; Hans H. Landsberg. Journal of the Royal Statistical Society, 107(1):64–65, ???? 1944. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2981357.

R:1944:BRBa

[319] F. E. R. Book review: Cost Behaviour and Price Policy. A Study Prepared, by the Committee on Price Determination for the Conference on Price Research. Journal of the Royal Statistical Society, 107(1):65–66, ???? 1944. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2981358.

R:1944:BRBb

[320] F. E. R. Book review: The Canadian Born in the United States: An Analysis of the Statistics of the Canadian Element in the Population of the United States, 1850 to 1930, by Leon E. Truesdell. Journal of the Royal Statistical Society, 107(1):66–67, ???? 1944. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981359.

Anonymous:1944:SNa

[321] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 107 (1):68–71, ???? 1944. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981360.

Anonymous:1944:CNa

[322] Anonymous. Current notes. Journal of the Royal Statistical Society, 107(1): 72–73, ???? 1944. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981361.

Yule:1944:RHH

[323] G. Udny Yule. Reginald Hawthorn Hooker, M.A. Journal of the Royal Statistical Society, 107(1):74–77, ???? 1944. ISSN 0952-8385. URL http://www.

jstor.org/stable/2981362.

Anonymous:1944:RHM

[324] Anonymous. Robert Holland-Martin, C.B. Journal of the Royal Statistical Society, 107(1):77–78, ???? 1944. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981363.

Anonymous:1944:SEAa

[325] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 107(1): 79–81, ???? 1944. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981364.

Anonymous:1944:LALa

[326] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 107(1):82–88, ???? 1944. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981365.

Anonymous:1944:PRa

[327] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 107 (1):89–91, ???? 1944. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981366.

Anonymous:1944:BMa

[328] Anonymous. Back matter. Journal of the Royal Statistical Society, 107(1):??, ???? 1944. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2981367.

Anonymous:1944:FMb

[329] Anonymous. Front matter. Journal of the Royal Statistical Society, 107(2):??, ???? 1944. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2981276.

Florence:1944:SIS

[330] P. Sargant Florence. The selection of industries suitable for dispersion into rural areas. Journal of the Royal Statistical Society, 107(2):93–116, ???? 1944. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2981277.

Pearce:1944:SMM

[331] S. C. Pearce. Sampling methods for the measurement of fruit crops. Journal of the Royal Statistical Society, 107(2): 117–126, ???? 1944. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981278.

P:1944:BRB

[332] E. S. P. Book review: The Advanced Theory of Statistics, by Maurice G. Kendall. Journal of the Royal Statistical Society, 107(2):127–129, ???? 1944. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2981279.

I:1944:BRB

[333] L. I. Book review: The Statistical Study of Literary Vocabulary, by G. Udny Yule. Journal of the Royal Statistical Society, 107(2):129–131, ???? 1944. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981280.

G:1944:BRBb

[334] I. G. G. Book review: The Problem of Valuation for Rating, by J. R. Hicks; U. K. Hicks; C. E. V. Leser. Journal of the Royal Statistical Society, 107(2): 131–135, ???? 1944. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981281.

Anonymous:1944:SNb

[335] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 107(2): 136–142, ???? 1944. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981282.

Anonymous:1944:CNb

[336] Anonymous. Current notes. Journal of the Royal Statistical Society, 107(2): 142–144, ???? 1944. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981283.

Anonymous:1944:SEAb

[337] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 107(2): 145–146, ???? 1944. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981284.

Anonymous:1944:PRb

[338] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 107(2): 147–150, ???? 1944. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981285.

Anonymous:1944:BMb

[339] Anonymous. Back matter. Journal of the Royal Statistical Society, 107(2):??, ???? 1944. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2981286.

Anonymous:1944:VI

[340] Anonymous. Volume information. Journal of the Royal Statistical Society, 107 (3/4):327–iv, ???? 1944. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2981211.

Anonymous:1944:FMc

[341] Anonymous. Front matter. Journal of the Royal Statistical Society, 107 (3/4):??, ???? 1944. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981212.

Box:1944:WSS

[342] Kathleen Box and Geoffrey Thomas. The wartime social survey. Journal of the Royal Statistical Society, 107(3/4): 151–189, ???? 1944. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981213.

Silvey:1944:MLR

[343] R. J. E. Silvey. Methods of listener research employed by the British Broadcasting Corporation. Journal of the Royal Statistical Society, 107(3/4):190– 230, ???? 1944. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981214.

Smith:1944:SDP

[344] J. L. Spencer Smith. The specification of disturbed periodic time series of the type of Wolfer’s annual sunspot numbers. Journal of the Royal Statistical Society, 107(3/4):231–241, ???? 1944.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2981215.

Richardson:1944:DWT

[345] Lewis F. Richardson. The distribution of wars in time. Journal of the Royal Statistical Society, 107(3/4):242–250, ???? 1944. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2981216. See

errata [367].

Geary:1944:CNV

[346] R. C. Geary. The concept of net volume of output, with special reference to Irish data. Journal of the Royal Statistical Society, 107(3/4):251–261, ???? 1944. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981217.

Anonymous:1944:WP

[347] Anonymous. Wholesale prices in 1943.

Journal of the Royal Statistical Society, 107(3/4):262–276, ???? 1944. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981218.

Anonymous:1944:RCF

[348] Anonymous. Report of the Council for the financial year ended December 31st, 1943, and for the sessional year ending June 20th, 1944, presented at the one hundred and tenth Annual General Meeting of the Royal Statistical Society, held at the London School of Hygiene, W.C. 1, on June 20th, 1944. Journal of the Royal Statistical Society, 107(3/4): 277–287, ???? 1944. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981219.

Anonymous:1944:POH

[349] Anonymous. Proceedings of the one hundred and tenth Annual General Meeting of the Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine on Tuesday, June 20th, 1944, at 5 p.m. Journal of the Royal Statistical Society, 107(3/4): 288–289, ???? 1944. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981220.

Geary:1944:RML

[350] R. C. Geary. Rejoinder to Messrs. Leak and Maizels. Journal of the Royal Statistical Society, 107(3/4):290–292, ???? 1944. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2981221. See

[247].

K:1944:BRBa

[351] M. G. K. Book review: Theory of Games and Economic Behaviour, by John von Neumann; Oskar Morgenstern. Journal of the Royal Statistical Society, 107 (3/4):293, ???? 1944. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981222.

K:1944:BRBb

[352] M. G. K. Book review: World Wheat Production: Its Regional Fluctuations and Inter-Regional Correlations, by V. P. Timoshenko. Journal of the Royal Statistical Society, 107(3/4):294– 295, ???? 1944. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981223.

G:1944:BRBc

[353] R. G. Book review: Economic Destiny, by R. G. Hawtrey. Journal of the Royal Statistical Society, 107(3/4): 295–296, ???? 1944. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981224.

R:1944:BRBc

[354] T. M. R. Book review: Full Employment in a Free Society, by William Beveridge. Journal of the Royal Statistical Society, 107(3/4):296–297, ???? 1944. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981225.

B:1944:BRB

[355] G. A. B. Book review: Foreign Exchange, by F. J. Docker. Journal of the Royal Statistical Society, 107(3/4): 298–299, ???? 1944. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981226.

R:1944:BRBd

[356] T. M. R. Book review: Comparative Operating Experience of Consumer Instalment Financing Agencies and Commercial Banks, 1929–41, by Ernst A. Dauer. Journal of the Royal Statistical Society, 107(3/4):299–300, ???? 1944. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981227.

R:1944:BRBe

[357] T. M. R. Book review: Future Education and Training for British Industry, by Tudor J. Jones. Journal of the Royal Statistical Society, 107(3/4):300, ???? 1944. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2981228.

G:1944:BRBd

[358] E. G. Book review: The Journey Home, by John Murray. Journal of the Royal Statistical Society, 107(3/4): 300–301, ???? 1944. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981229.

Anonymous:1944:SNc

[359] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 107 (3/4):302–307, ???? 1944. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2981230.

Anonymous:1944:CNc

[360] Anonymous. Current notes. Journal of the Royal Statistical Society, 107 (3/4):308, ???? 1944. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981231.

Anonymous:1944:SEAc

[361] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 107(3/4): 309–311, ???? 1944. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981232.

Anonymous:1944:LALb

[362] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 107(3/4):312–320, ???? 1944. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2981233.

Anonymous:1944:PRL

[363] Anonymous. Periodicals received by the Library. Journal of the Royal Statistical Society, 107(3/4):321–326, ???? 1944. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2981234.

Anonymous:1944:BMc

[364] Anonymous. Back matter. Journal of the Royal Statistical Society, 107 (3/4):??, ???? 1944. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981235.

Anonymous:1945:VIa

[365] Anonymous. Volume information. Journal of the Royal Statistical Society, 108 (1/2):iii–iv, ???? 1945. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981189.

Anonymous:1945:FMa

[366] Anonymous. Front matter. Journal of the Royal Statistical Society, 108(1/2): i–ii, ???? 1945. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981190.

Richardson:1945:EDW

[367] L. F. Richardson. Errata: The distribution of wars in time. Journal of the Royal Statistical Society, 108 (1/2):??, ???? 1945. ISSN 0952-8385.

URL http://www.jstor.org/stable/

2981191. See [345].

Elderton:1945:CSS

[368] William Elderton and George H. Wood.

Cricket scores and some skew correlation distributions: (an arithmetical study). Journal of the Royal Statistical Society, 108(1/2):1–11, ???? 1945. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981192.

Elderton:1945:CSG

[369] William Elderton and George H. Wood.

Cricket scores and geometrical progression. Journal of the Royal Statistical Society, 108(1/2):12–40, ???? 1945. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981193. See errata [389].

Sutton:1945:EAS

[370] Charles T. Sutton. An enquiry into the assessing of salvage awards. Journal of the Royal Statistical Society, 108(1/2): 41–92, ???? 1945. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981194.

Kendall:1945:AOT

[371] M. G. Kendall. On the analysis of oscillatory time-series. Journal of the Royal Statistical Society, 108(1/2):93– 141, ???? 1945. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981195.

Leak:1945:SBI

[372] H. Leak and A. Maizels. The structure of British industry. Journal of the Royal Statistical Society, 108(1/2): 142–207, ???? 1945. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981196.

Yule:1945:MST

[373] G. Udny Yule. On a method of studying time-series based on their internal correlations. Journal of the Royal Statistical Society, 108(1/2):208–225, ???? 1945. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981197.

Kendall:1945:NMY

[374] M. G. Kendall. Note on Mr. Yule’s paper. Journal of the Royal Statistical Society, 108(1/2):226–230, ???? 1945. ISSN 0952-8385. URL http://www.

jstor.org/stable/2981198.

Anonymous:1945:RRI

[375] Anonymous. Reconstitution of the research and industrial applications sections. Journal of the Royal Statistical Society, 108(1/2):231–234, ???? 1945. ISSN 0952-8385. URL http://www.

jstor.org/stable/2981199.

Anonymous:1945:BRB

[376] Anonymous. Book review: Statistical Method for Research Workers, by R. A. Fisher. Journal of the Royal Statistical Society, 108(1/2):235, ???? 1945. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981200.

R:1945:BRBa

[377] F. E. R. Book review: The Mining Industries, 1899–1939: A Study of Output, Employment and Productivity, by Harold Barger; Sam H. Schurr. Journal of the Royal Statistical Society, 108 (1/2):235–236, ???? 1945. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2981201.

R:1945:BRBb

[378] F. E. R. Book review: Introduction to Economics, by Alec Cairncross. Journal of the Royal Statistical Society, 108 (1/2):236, ???? 1945. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981202.

R:1945:BRBc

[379] T. M. R. Book review: Corporate Cash Balances, 1914–1943, by Friedrich A. Lutz. Journal of the Royal Statistical Society, 108(1/2):237–238, ???? 1945. ISSN 0952-8385. URL http://www.

jstor.org/stable/2981203.

B:1945:BRB

[380] G. A. B. Book review: Accounts Analysis, by J. A. Mackinnon; L. J. E. Troupe; Design of Accounts. by F. Sewell Bray; H. Basil Sheasby; Balance Sheet Values. by P. D. Leake. Journal of the Royal Statistical Society, 108(1/2):238–239, ???? 1945. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981204.

Anonymous:1945:SNa

[381] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 108 (1/2):240–243, ???? 1945. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2981205.

Anonymous:1945:CNa

[382] Anonymous. Current notes. Journal of the Royal Statistical Society, 108(1/2): 243–244, ???? 1945. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981206.

H:1945:SDW

[383] A. B. H. Sir Duncan Wilson, C.B.E., C.V.O. Journal of the Royal Statistical Society, 108(1/2):245, ???? 1945. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2981207.

Anonymous:1945:SEAa

[384] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 108(1/2): 246–250, ???? 1945. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981208.

Anonymous:1945:LALa

[385] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 108(1/2):251–255, ???? 1945. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2981209.

Anonymous:1945:PRa

[386] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 108 (1/2):256–258, ???? 1945. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2981210.

Anonymous:1945:VIb

[387] Anonymous. Volume information. Journal of the Royal Statistical Society, 108 (3/4):512–514, ???? 1945. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2981287.

Anonymous:1945:FMb

[388] Anonymous. Front matter. Journal of the Royal Statistical Society, 108(3/4): i–iii, ???? 1945. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981288.

Anonymous:1945:ECS

[389] Anonymous. Errata: Cricket scores.

Journal of the Royal Statistical Society, 108(3/4):??, ???? 1945. ISSN 0952-8385.

URL http://www.jstor.org/stable/

2981289. See [369].

Butler:1945:FRM

[390] William Butler. The fatality rate of measles: A study of its trend in time. Journal of the Royal Statistical Society, 108(3/4):259–285, ???? 1945. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981290.

Stone:1945:AMD

[391] Richard Stone. The analysis of market demand. Journal of the Royal Statistical Society, 108(3/4):286–391, ???? 1945. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981291.

Frankel:1945:IDP

[392] H. Frankel. The industrial distribution of the population of Great Britain in July, 1939. Journal of the Royal Statistical Society, 108(3/4):392–430, ???? 1945. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981292.

Anonymous:1945:WP

[393] Anonymous. Wholesale prices in 1944. Journal of the Royal Statistical Society, 108(3/4):431–446, ???? 1945. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981293.

Anonymous:1945:MDN

[394] Anonymous. The mathematics division of the national physical laboratory. Journal of the Royal Statistical Society, 108(3/4):447–448, ???? 1945. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981294.

Anonymous:1945:RCF

[395] Anonymous. Report of the Council for the financial year ended December 31st, 1944, and for the sessional year ending June 26th, 1945, presented at the one hundred and eleventh Annual General Meeting of the Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine, W.C. 1, on June 26th, 1945. Journal of the Royal Statistical Society, 108(3/4):449–461, ???? 1945. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981295.

Anonymous:1945:POH

[396] Anonymous. Proceedings of the one hundred and eleventh Annual General Meeting of the Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine on Tuesday, June 26th, 1945, at 5 p.m. Journal of the Royal Statistical Society, 108(3/4): 462–465, ???? 1945. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981296.

R:1945:BRBd

[397] E. C. R. Book review: The Adolescent Criminal, by W. Norwood East; Percy Stocks; H. T. P. Young. Journal of the Royal Statistical Society, 108(3/4): 466–467, ???? 1945. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981297.

R:1945:BRBe

[398] H. W. R. Book review: Outlay and Income in the United States 1921–38, by Harold Barger. Journal of the Royal Statistical Society, 108(3/4):467–468, ???? 1945. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981298.

R:1945:BRBf

[399] H. W. R. Book review: The Economics of Full Employment, by Oxford University Institute of Statistics. Journal of the Royal Statistical Society, 108(3/4): 468–470, ???? 1945. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981299.

R:1945:BRBg

[400] T. M. R. Book review: Full Employment and Free Trade, by Michael Polanyi. Journal of the Royal Statistical Society, 108(3/4):470–471, ???? 1945. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981300.

R:1945:BRBh

[401] T. M. R. Book review: The Pattern of Corporate Financial Structure, by Walter A. Chudson. Journal of the Royal Statistical Society, 108(3/4):472, ???? 1945. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981301.

Anonymous:1945:SNb

[402] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 108 (3/4):473–477, ???? 1945. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2981302.

Anonymous:1945:CNb

[403] Anonymous. Current notes. Journal of the Royal Statistical Society, 108(3/4): 477–478, ???? 1945. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981303.

L:1945:SPA

[404] H. L. Sir Percy Ashley, K.B.E., C.B. Journal of the Royal Statistical Society, 108(3/4):479–480, ???? 1945. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2981304.

Bowley:1945:SHL

[405] A. L. Bowley. Sir Hubert Llewellyn Smith, G.C.B. Journal of the Royal Statistical Society, 108(3/4):480–481, ???? 1945. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2981305.

Collet:1945:CBD

[406] Clara E. Collet. Charles Booth, the Dension Club and H. Llewellyn Smith. Journal of the Royal Statistical Society, 108(3/4):482–485, ???? 1945. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2981306.

Bowley:1945:GHW

[407] A. L. Bowley and E. C. R. George Henry wood. Journal of the Royal Statistical Society, 108(3/4):485–487, ???? 1945. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2981307.

R:1945:JFT

[408] G. R. James Fowler Tocher, 1864–1945.

Journal of the Royal Statistical Society, 108(3/4):487–489, ???? 1945. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981308.

Anonymous:1945:SEAb

[409] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 108(3/4): 490–493, ???? 1945. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981309.

Anonymous:1945:LALb

[410] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 108(3/4):494–500, ???? 1945. ISSN 0952-8385. URL http://www.

jstor.org/stable/2981310.

Anonymous:1945:PRL

[411] Anonymous. Periodicals received by the Library. Journal of the Royal Statistical Society, 108(3/4):501–507, ???? 1945. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981311.

Anonymous:1945:PRb

[412] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 108 (3/4):508–511, ???? 1945. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2981312.

Anonymous:1946:VIa

[413] Anonymous. Volume information. Journal of the Royal Statistical Society, 109 (1):iii–iv, ???? 1946. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981387.

Anonymous:1946:FM

[414] Anonymous. Front matter. Journal of the Royal Statistical Society, 109(1):i–ii, ???? 1946. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2981388.

Anonymous:1946:PFO

[415] Anonymous. Proceedings of the first ordinary meeting of the session, 1945– 46, held on November 22nd, 1945, the President, the Rt. Hon. Lord Woolton, P.C., C.H., in the Chair. Journal of the Royal Statistical Society, 109(1):1–11, ???? 1946. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2981389.

Yates:1946:RRS

[416] F. Yates. A review of recent statistical developments in sampling and sampling surveys. Journal of the Royal Statistical Society, 109(1):12–43, ???? 1946. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981390.

Yule:1946:CSS

[417] G. Udny Yule. Cumulative sampling: A speculation as to what happens in copying manuscripts. Journal of the Royal Statistical Society, 109(1):44–52, ???? 1946. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2981391.

Penrose:1946:ESM

[418] L. S. Penrose. The elementary statistics of majority voting. Journal of the Royal Statistical Society, 109(1):53–57, ???? 1946. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2981392.

Anonymous:1946:RRI

[419] Anonymous. Regulations of the research and industrial applications sections. Journal of the Royal Statistical Society, 109(1):58–60, ???? 1946. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981393.

R:1946:BRBa

[420] T. M. R. Book review: Commercial Policy in the Post-War World. A Report, by the Economic and Financial Committees of the League of Nations. Journal of the Royal Statistical Society, 109(1):61, ???? 1946. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981394.

Z:1946:BRB

[421] A. P. Z. Book review: The Exploitation of Foreign Labour, by Germany. Journal of the Royal Statistical Society, 109 (1):61–62, ???? 1946. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981395.

L:1946:BRB

[422] A. M. L. Book review: Radio Advertising for Retailers, by C. H. Sandage. Journal of the Royal Statistical Society, 109(1):62–63, ???? 1946. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2981396.

Anonymous:1946:SNa

[423] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 109 (1):64–68, ???? 1946. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981397.

Anonymous:1946:CNa

[424] Anonymous. Current notes. Journal of the Royal Statistical Society, 109(1):69, ???? 1946. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2981398.

Woolton:1946:MER

[425] Lord Woolton. Miss Eleanor Rathbone, M.P. Journal of the Royal Statistical Society, 109(1):70, ???? 1946. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2981399.

Anonymous:1946:SEAa

[426] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 109(1): 71–74, ???? 1946. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981400.

Anonymous:1946:LALa

[427] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 109(1):75–77, ???? 1946. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2981401.

Anonymous:1946:PRa

[428] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 109 (1):78–83, ???? 1946. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981402.

Greenwood:1946:SSI

[429] Major Greenwood. The statistical study of infectious diseases. Journal of the Royal Statistical Society, 109(2):85–110, ???? 1946. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2981176.

Mandeville:1946:IMC

[430] John P. Mandeville. Improvements in methods of census and survey analysis. Journal of the Royal Statistical Society, 109(2):111–129, ???? 1946. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2981177.

Richardson:1946:NNE

[431] Lewis F. Richardson. The number of nations on each side of a war. Journal of the Royal Statistical Society, 109(2): 130–156, ???? 1946. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981178.

Irwin:1946:IWB

[432] J. O. Irwin. On the interpretation of within and between class analysis of variance when the intra class correlation is negative. Journal of the Royal Statistical Society, 109(2):157–158, ???? 1946. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981179.

K:1946:BRBa

[433] O. K. Book review: Industrial Experimentation, by K. A. Brownlee. Journal of the Royal Statistical Society, 109 (2):159, ???? 1946. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981180.

K:1946:BRBb

[434] J. K. Book review: Intergovernmental Commodity Control Agreements. ; Commodity Control: A Study of Primary Products, by P. Lamartine Yates. Journal of the Royal Statistical Society, 109 (2):159–160, ???? 1946. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2981181.

S:1946:BRBa

[435] H. W. S. Book review: A Short History of Labour Conditions under Industrial Capitalism. Vol. 3,Part 1. Germany, 1800 to the Present Day, by Jurgen Kuczynski. Journal of the Royal Statistical Society, 109(2):160–161, ????

1946. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2981182.

Anonymous:1946:SNb

[436] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 109(2): 162–166, ???? 1946. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981183.

Anonymous:1946:CNb

[437] Anonymous. Current notes. Journal of the Royal Statistical Society, 109(2): 166–168, ???? 1946. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981184.

H:1946:LK

[438] R. G. H. Lord Keynes. Journal of the Royal Statistical Society, 109(2):169, ???? 1946. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2981185.

L:1946:SAC

[439] H. L. Sedley Anthony Cudmore. Journal of the Royal Statistical Society, 109(2): 169–170, ???? 1946. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981186.

Anonymous:1946:SEAb

[440] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 109(2): 171–173, ???? 1946. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981187.

Anonymous:1946:LALb

[441] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 109(2):174–177, ???? 1946. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2981188.

Blau:1946:WWE

[442] Gerda Blau. Wool in the world economy. Journal of the Royal Statistical Society, 109(3):179–242, ???? 1946. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981368.

Allen:1946:MAB

[443] R. G. D. Allen. Mutual aid between the U.S. and the British Empire, 1941– 45. Journal of the Royal Statistical Society, 109(3):243–277, ???? 1946. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981369.

Shaul:1946:DTF

[444] J. R. H. Shaul. Derivation of total fertility, Gross and net reproduction rates from census statistics of marriage fertility. Journal of the Royal Statistical Society, 109(3):278–283, ???? 1946. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981370.

Kempthorne:1946:UPC

[445] O. Kempthorne. The use of a punchedcard system for the analysis of survey data, with special reference to the analysis of the national farm survey. Journal of the Royal Statistical Society, 109(3): 284–295, ???? 1946. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981371.

Skellam:1946:FDD

[446] J. G. Skellam. The frequency distribution of the difference between two Poisson variates belonging to different populations. Journal of the Royal Statistical Society, 109(3):296, ???? 1946. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981372.

A:1946:BRBa

[447] R. G. D. A. Book review: Rudimentary Mathematics for Economists and Statisticians, by W. L. Crum; Joseph A. Schumpeter. Journal of the Royal Statistical Society, 109(3):297, ???? 1946. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2981373.

S:1946:BRBb

[448] A. W. S. Book review: Sampling Inspection Tables: Single and Doubling Sampling, by H. F. Dodge; H. G. Romig. Journal of the Royal Statistical Society, 109(3):297–298, ???? 1946. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2981374.

K:1946:BRBc

[449] M. G. K. Book review: Measuring Business Cycles, by Arthur F. Burns; Wesley C. Mitchell. Journal of the Royal Statistical Society, 109(3):298–299, ???? 1946. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2981375.

R:1946:BRBb

[450] T. M. R. Book review: Changes in Income Distribution During the Great Depression, by Horst Mendershausen. Journal of the Royal Statistical Society, 109(3):299–300, ???? 1946. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2981376.

B:1946:BRBa

[451] T. B. Book review: Price Flexibility and Employment, by Oscar Lange. Journal of the Royal Statistical Society, 109(3): 300–301, ???? 1946. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981377.

B:1946:BRBb

[452] G. A. B. Book review: Price Control and Business, by George Katona. Journal of the Royal Statistical Society, 109(3): 301–302, ???? 1946. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981378.

Anonymous:1946:ONP

[453] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 109(3):302–303, ???? 1946. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981379.

Anonymous:1946:SNc

[454] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 109(3): 304–307, ???? 1946. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981380.

Anonymous:1946:CNc

[455] Anonymous. Current notes. Journal of the Royal Statistical Society, 109(3): 308–310, ???? 1946. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981381.

M:1946:SPH

[456] F. A. A. M. Sir Percy Harvey, K.B.E., C.B., F.I.A. Journal of the Royal Statistical Society, 109(3):311–312, ???? 1946. ISSN 0952-8385. URL http://www.

jstor.org/stable/2981382.

H:1946:GRD

[457] A. B. H. George Rae, D.Sc. 1884–1946.

Journal of the Royal Statistical Society, 109(3):313, ???? 1946. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981383.

Anonymous:1946:SEAc

[458] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 109(3): 314–316, ???? 1946. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981384.

Anonymous:1946:LALc

[459] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 109(3):317–319, ???? 1946. ISSN 0952-8385. URL http://www.

jstor.org/stable/2981385.

Anonymous:1946:PRb

[460] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 109(3): 320–323, ???? 1946. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981386.

Anonymous:1946:VIb

[461] Anonymous. Volume information. Journal of the Royal Statistical Society, 109 (4):523–525, ???? 1946. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2981329.

Greenwood:1946:PMR

[462] M. Greenwood. Proceedings of a meeting of the Royal Statistical Society held on July 16th, 1946. Journal of the Royal Statistical Society, 109(4):325–378, ???? 1946. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981330.

Anonymous:1946:WP

[463] Anonymous. Wholesale prices in 1945. Journal of the Royal Statistical Society, 109(4):379–394, ???? 1946. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981331.

Rigg:1946:RAM

[464] F. A. Rigg. Recent advances in mathematical statistics. Journal of the Royal Statistical Society, 109(4):395–450, ???? 1946. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981332.

Busk:1946:DSD

[465] Th. Busk. Dot and scatter diagrams made by means of a typewriter. Journal of the Royal Statistical Society, 109(4): 451–456, ???? 1946. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981333.

Plackett:1946:LTE

[466] R. L. Plackett. Literature on testing the equality of variances and covariances in normal populations. Journal of the Royal Statistical Society, 109(4): 457–468, ???? 1946. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981334.

Kempthorne:1946:SCC

[467] O. Kempthorne and D. A. Boyd. The stock-carrying capacity of farms. Journal of the Royal Statistical Society, 109 (4):469–475, ???? 1946. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2981335.

Anonymous:1946:RCF

[468] Anonymous. Report of the Council for the financial year ended December 31st, 1945, and for the sessional year ending June 27th, 1946, presented at the one hundred and twelfth Annual General Meeting of the Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine, W.C. 1, on June 27th, 1946. Journal of the Royal Statistical Society, 109(4):476–489, ????

1946. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981336.

Anonymous:1946:POH

[469] Anonymous. Proceedings of the one hundred and twelfth Annual General Meeting of the Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine on Thursday, June 27th, 1946, at 5 p.m. Journal of the Royal Statistical Society, 109(4): 490–491, ???? 1946. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981337.

Anonymous:1946:MPL

[470] Anonymous. Memorandum on the proposal laid before the special general meeting of 7th November, 1946, that the Society’s petition for a supplemental charter be withdrawn. Journal of the Royal Statistical Society, 109(4):492– 502, ???? 1946. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981338.

V:1946:BRB

[471] S. V. Book review: Fundamentals of the Theory of Statistics.: Vol. 1. Elementary Statistics and Applications, by J. G. Smith; A. J. Duncan; Fundamentals of the Theory of Statistics.: Vol.

2. Sampling Statistics and Applications. by J. G. Smith; A. J. Duncan. Journal of the Royal Statistical Society, 109(4): 503–504, ???? 1946. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981339.

A:1946:BRBb

[472] F. J. A. Book review: Contributions to the Study of Oscillatory Time-Series, by M. G. Kendall. Journal of the Royal Statistical Society, 109(4):504, ???? 1946.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2981340.

A:1946:BRBc

[473] F. J. A. Book review: Sequential Analysis of Statistical Data: Applications. Journal of the Royal Statistical Society, 109(4):505, ???? 1946. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981341.

W:1946:BRBa

[474] J. W. Book review: Statistical Methods Applied to Experiments in Agriculture and Biology, by George W. Snedecor. Journal of the Royal Statistical Society, 109(4):505–506, ???? 1946. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981342.

G:1946:BRBa

[475] M. G. Book review: Statistiche Methoden fur Naturwissenschafter, Mediziner und Ingenieure, by Arthur Linder. Journal of the Royal Statistical Society, 109 (4):506, ???? 1946. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981343.

W:1946:BRBb

[476] J. W. Book review: A First Course in Mathematical Statistics, by C. E. Weatherburn. Journal of the Royal Statistical Society, 109(4):506–507, ???? 1946. ISSN 0952-8385. URL http://www.

jstor.org/stable/2981344.

G:1946:BRBb

[477] R. G. Book review: International Implications of Full Employment in Great Britain, by Allan G. B. Fisher. Journal of the Royal Statistical Society, 109(4): 508–509, ???? 1946. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981345.

C:1946:BRB

[478] R. D. C. Book review: The Money Value of a Man, by Louis I. Dublin; Alfred J. Lotka. Journal of the Royal Statistical Society, 109(4):509, ???? 1946. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2981346.

E:1946:BRB

[479] W. P. E. Book review: A Century of American Life Insurance. A History of the Mutual Life Insurance Company of New York 1843–1943, by Shepard B. Clough. Journal of the Royal Statistical Society, 109(4):509–510, ???? 1946. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2981347.

Anonymous:1946:SNd

[480] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 109(4): 511–515, ???? 1946. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981348.

Anonymous:1946:SEAd

[481] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 109(4): 516–518, ???? 1946. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981349.

Anonymous:1946:LALd

[482] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 109(4):519–522, ???? 1946. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2981350.

Anonymous:1947:VIa

[483] Anonymous. Volume information. Journal of the Royal Statistical Society, 110 (1):??, ???? 1947. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981258.

Anonymous:1947:FM

[484] Anonymous. Front matter. Journal of the Royal Statistical Society, 110(1):??, ???? 1947. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2981259.

Anonymous:1947:PFO

[485] Anonymous. Proceedings of the first ordinary meeting of the session, 1946–47, held on November 27th, 1946, the President, the Rt. Hon. Lord Woolton, P. C., C. H., in the Chair. Journal of the Royal Statistical Society, 110(1):1–26, ???? 1947. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2981260.

Wishart:1947:SAD

[486] John Wishart. Statistical aspects of demobilization in the Royal Navy. Journal of the Royal Statistical Society, 110 (1):27–50, ???? 1947. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981261.

Anonymous:1947:RTS

[487] Anonymous. Report on the teaching of statistics in universities and university colleges. Journal of the Royal Statistical Society, 110(1):51–57, ???? 1947. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981262.

Anonymous:1947:FRE

[488] Anonymous. First report of the Examinations Committee, as amended and adopted by the Council. Journal of the Royal Statistical Society, 110(1):58–71, ???? 1947. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2981263.

C:1947:BRBa

[489] D. G. C. Book review: Time Series Analysis: Smoothing, by Stages. by Lewis A. Maverick. Journal of the Royal Statistical Society, 110(1):72–74, ???? 1947. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2981264.

V:1947:BRBa

[490] S. V. Book review: An Experimental Introduction to the Theory of Probability, by J. E. Kerrich. Journal of the Royal Statistical Society, 110(1):74, ???? 1947. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981265.

C:1947:BRBb

[491] E. A. C. Book review: The Registrar-General’s Statistical Review of England and Wales for the Years 1938 and 1939. Journal of the Royal Statistical Society, 110(1):75–77, ???? 1947. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2981266.

R:1947:BRBa

[492] T. M. R. Book review: Income from Independent Professional Practice, by Milton Friedman; Simon Kuznets. Journal of the Royal Statistical Society, 110 (1):77–78, ???? 1947. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981267.

F:1947:BRB

[493] P. S. F. Book review: Reconversion, Modernization and Expansion. Journal of the Royal Statistical Society, 110 (1):78, ???? 1947. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981268.

D:1947:BRB

[494] W. M. D. Book review: Bretton Woods, for Better or Worse, by R. G. Hawtrey. Journal of the Royal Statistical Society, 110(1):78–79, ???? 1947. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2981269.

Anonymous:1947:SNa

[495] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 110 (1):80–82, ???? 1947. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981270.

Anonymous:1947:CNa

[496] Anonymous. Current notes. Journal of the Royal Statistical Society, 110(1): 83–84, ???? 1947. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981271.

H:1947:IF

[497] R. G. H. Irving Fisher. Journal of the Royal Statistical Society, 110(1):85, ???? 1947. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981272.

N:1947:MH

[498] J. W. N. Michel Huber. Journal of the Royal Statistical Society, 110(1):85–86, ???? 1947. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2981273.

Anonymous:1947:SEAa

[499] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 110(1): 87–90, ???? 1947. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981274.

Anonymous:1947:PRa

[500] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 110 (1):91–94, ???? 1947. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981275.

Weeks:1947:SSI

[501] H. T. Weeks. Statistics and the statistician in industry. Journal of the Royal Statistical Society, 110(2):95–107, ???? 1947. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2981313.

Tippett:1947:SIE

[502] L. H. C. Tippett. The study of industrial efficiency, with special reference to the cotton industry. Journal of the Royal Statistical Society, 110(2):108–122, ???? 1947. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2981314.

Quenouille:1947:LST

[503] M. H. Quenouille. A large-sample test for the goodness of fit of autoregressive schemes. Journal of the Royal Statistical Society, 110(2):123–129, ???? 1947. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2981315.

Kendall:1947:BRB

[504] David G. Kendall. Book review: On Random Processes and their Application to Sickness and Accident Statistics. Journal of the Royal Statistical Society, 110(2):130–137, ???? 1947. URL http:

                                        //www.jstor.org/stable/2981316.

Percival:1947:ECT

[505] D. A. Percival. An experiment in census tabulation. Census of Cyprus, 1946. Journal of the Royal Statistical Society, 110(2):138–143, ???? 1947. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981317.

P:1947:BRB

[506] E. S. P. Book review: The Advanced Theory of Statistics, by Maurice G. Kendall. Journal of the Royal Statistical Society, 110(2):144–145, ???? 1947. ISSN 0952-8385. URL http://www.

jstor.org/stable/2981318.

I:1947:BRBa

[507] J. O. I. Book review: Introduction to Mathematical Statistics, by Paul G. Hoel. Journal of the Royal Statistical Society, 110(2):145–146, ???? 1947. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981319.

V:1947:BRBb

[508] S. V. Book review: Statistical Quality Control, by Eugene L. Grant. Journal of the Royal Statistical Society, 110(2): 146–147, ???? 1947. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981320.

W:1947:BRB

[509] L. W. Book review: A Chapter in Population Sampling. Journal of the Royal Statistical Society, 110(2):147, ???? 1947. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2981321.

H:1947:BRBa

[510] R. G. H. Book review: A History of Banking Theory in Great Britain and the United States, by Lloyd W. Mints. Journal of the Royal Statistical Society, 110(2):147–148, ???? 1947. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981322.

G:1947:BRBa

[511] M. G. Book review: John Hilton: The Story of His Life, by Edna Nixon. Journal of the Royal Statistical Society, 110 (2):148–149, ???? 1947. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2981323.

Anonymous:1947:SNb

[512] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 110(2): 150–153, ???? 1947. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981324.

Anonymous:1947:CNb

[513] Anonymous. Current notes. Journal of the Royal Statistical Society, 110(2):154, ???? 1947. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2981325.

Anonymous:1947:SEAb

[514] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 110(2): 155–158, ???? 1947. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981326.

Anonymous:1947:LALa

[515] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 110(2):159–164, ???? 1947. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2981327.

Anonymous:1947:PRb

[516] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 110(2): 165–168, ???? 1947. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981328.

Snow:1947:SBE

[517] E. C. Snow. The statistical basis of export targets. Journal of the Royal Statistical Society, 110(3):169–186, ???? 1947. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981403.

Marley:1947:SES

[518] Joan G. Marley. A statistical and economic survey of certain aspects of the beef producing, milk producing and cattle rearing industries in Great Britain between 1939 and 1945. Journal of the Royal Statistical Society, 110(3):187– 256, ???? 1947. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981404.

Smith:1947:SEM

[519] H. Fairfield Smith. Standard errors of means in sampling surveys with twostage sampling. Journal of the Royal Statistical Society, 110(3):257–259, ???? 1947. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981405.

Anonymous:1947:RCNa

[520] Anonymous. Reports from Council and notices to Fellows. Journal of the Royal Statistical Society, 110(3):260–262, ???? 1947. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981406.

I:1947:BRBb

[521] J. O. I. Book review: Probit Analysis: A Statistical Treatment of the Sigmoid Response Curve, by D. J. Finney; F. Tattersfield. Journal of the Royal Statistical Society, 110(3):263–266, ???? 1947. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981407.

V:1947:BRBc

[522] S. V. Book review: Statistical Methods in Research and Production, with Special Reference to the Chemical Industry, by O. L. Davies. Journal of the Royal Statistical Society, 110(3):266–267, ???? 1947. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2981408.

E:1947:BRB

[523] W. P. E. Book review: Teorica della Sopravvivenza, by Filadelfo Insolera. Journal of the Royal Statistical Society, 110 (3):267, ???? 1947. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981409.

C:1947:BRBc

[524] M. J. R. C. Book review: Manual de Estatistica, by Amaro D. Guerreiro. Journal of the Royal Statistical Society, 110 (3):268, ???? 1947. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981410.

K:1947:BRBa

[525] W. F. F. K. Book review: Food, Famine and Relief, 1940–46. Journal of the Royal Statistical Society, 110(3): 268–269, ???? 1947. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981411.

B:1947:BRB

[526] W. B. Book review: The Population and Epidemics of Exeter in Pre-Census Times, by Ransom Pickard; James Townsend. Journal of the Royal Statistical Society, 110(3):269–270, ???? 1947. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2981412.

C:1947:BRBd

[527] P. R. C. Book review: The Population of the Soviet Union: History and Prospects, by Frank Lorimer. Journal of the Royal Statistical Society, 110(3):270, ???? 1947. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2981413.

K:1947:BRBb

[528] J. K. Book review: Income: An Introduction to Economics, by A. C. Pigou. Journal of the Royal Statistical Society, 110(3):271–272, ???? 1947. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981414.

M:1947:BRB

[529] A. M. Book review: Small and Big Business: Economic Problems of the Size of Firm, by J. Steindl. Journal of the Royal Statistical Society, 110(3):272–274, ???? 1947. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981415.

Anonymous:1947:SNc

[530] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 110(3): 275–276, ???? 1947. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981416.

Anonymous:1947:CNc

[531] Anonymous. Current notes. Journal of the Royal Statistical Society, 110(3):277, ???? 1947. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2981417.

Anonymous:1947:SEAc

[532] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 110(3): 278–282, ???? 1947. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981418.

Anonymous:1947:VIb

[533] Anonymous. Volume information. Journal of the Royal Statistical Society, 110 (4):391–393, ???? 1947. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2981236.

Shone:1947:ISD

[534] Robert M. Shone. The iron and steel development plan: Some statistical considerations. Journal of the Royal Statistical Society, 110(4):283–309, ???? 1947. ISSN 0952-8385. URL http://www.

jstor.org/stable/2981237.

Kyd:1947:PTS

[535] J. G. Kyd. The proposed third statistical account of Scotland. Journal of the Royal Statistical Society, 110(4): 310–321, ???? 1947. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981238.

Anonymous:1947:WP

[536] Anonymous. Wholesale prices in 1946. Journal of the Royal Statistical Society, 110(4):322–336, ???? 1947. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981239.

Sichel:1947:FGF

[537] Herbert S. Sichel. Fitting growth and frequency curves by the method of frequency moments. Journal of the Royal Statistical Society, 110(4):337–347, ???? 1947. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981240.

Anonymous:1947:SSI

[538] Anonymous. Statistical and Social Inquiry Society of Ireland: Centenary celebrations. Journal of the Royal Statistical Society, 110(4):348, ???? 1947. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981241.

Geary:1947:SRT

[539] R. C. Geary. Some reflections on tendencies in the theory and practice of statistics in Ireland and elsewhere. Journal of the Royal Statistical Society, 110(4): 348–349, ???? 1947. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981242.

Anonymous:1947:RCF

[540] Anonymous. Report of the Council for the financial year ended December 31st, 1946, and for the sessional year ending July 10th, 1947, presented at the one hundred and thirteenth Annual General Meeting of the Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine, W.C. 1, on July 10th, 1947. Journal of the Royal Statistical Society, 110(4):350–366, ???? 1947. ISSN 0952-8385. URL http://

www.jstor.org/stable/2981243.

Anonymous:1947:RCNb

[541] Anonymous. Reports from Council and notices to Fellows. Journal of the Royal Statistical Society, 110(4):366, ???? 1947. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2981244.

V:1947:BRBd

[542] S. V. Book review: Selected Techniques of Statistical Analysis for Scientific and Industrial Research and Production and Management Engineering. Journal of the Royal Statistical Society, 110(4):367–368, ???? 1947. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981245.

N:1947:BRB

[543] D. N. Book review: Industrial Statistics and Quality Control, by Paul Peach. Journal of the Royal Statistical Society, 110(4):368–369, ???? 1947. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2981246.

H:1947:BRBb

[544] H. O. H. Book review: An Introduction to Managerial Business Statistics, by Harry Pelle Hartkemeier. Journal of the Royal Statistical Society, 110(4):369, ???? 1947. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2981247.

G:1947:BRBb

[545] R. F. G. Book review: Cost of Living Statistics: Methods and Techniques for the Post War Period. Journal of the Royal Statistical Society, 110(4):369– 370, ???? 1947. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2981248.

G:1947:BRBc

[546] R. F. G. Book review: Report of the President’s Committee on the Cost of Living. Journal of the Royal Statistical Society, 110(4):370–371, ???? 1947. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2981249.

R:1947:BRBb

[547] T. M. R. Book review: World Fiber Survey. Journal of the Royal Statistical Society, 110(4):371–372, ???? 1947. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2981250.

G:1947:BRBd

[548] R. F. G. Book review: Housing and the Family, by M. J. Elsas. Journal of the Royal Statistical Society, 110(4): 372–373, ???? 1947. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981251.

Anonymous:1947:SNd

[549] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 110(4): 374–378, ???? 1947. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981252.

Anonymous:1947:CNd

[550] Anonymous. Current notes. Journal of the Royal Statistical Society, 110(4): 379–381, ???? 1947. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981253.

G:1947:SIG

[551] R. F. G. Sir Ioan Gwilym Gibbon. Journal of the Royal Statistical Society, 110 (4):382–383, ???? 1947. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2981254.

G:1947:RRK

[552] D. V. G. Robert Rene Kuczynski, 1876– 1947. Journal of the Royal Statistical Society, 110(4):383–384, ???? 1947. ISSN

0952-8385. URL http://www.jstor.
org/stable/2981255.

Anonymous:1947:SEAd

[553] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 110(4): 385–387, ???? 1947. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2981256.

Anonymous:1947:LALb

[554] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 110(4):388–390, ???? 1947. ISSN 0952-8385. URL http://www.

jstor.org/stable/2981257.

Anonymous:1948:FMa

[555] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 111(1):??, ???? 1948.

URL http://www.jstor.org/stable/
2980719.

Heron:1948:FPS

[556] David Heron. On the financial position of the Society. Journal of the Royal Statistical Society. Series A (General), 111 (1):1–9, ???? 1948. URL http://www.

jstor.org/stable/2980720.

Anonymous:1948:PM

[557] Anonymous. Proceedings of the meeting. Journal of the Royal Statistical Society. Series A (General), 111(1):9–11,

???? 1948. URL http://www.jstor.
org/stable/2980721.

Anonymous:1948:PSG

[558] Anonymous. Proceedings of a special general meeting of the Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine on Thursday, 20th November, 1947, immediately on conclusion of the ordinary meeting convened for the same date. Journal of the Royal Statistical Society. Series A (General), 111(1):12–13,

???? 1948. URL http://www.jstor.
org/stable/2980722.

Cotton:1948:CMD

[559] H. Cotton and D. Heron. The collection of morbidity data from hospitals. Journal of the Royal Statistical Society. Series A (General), 111(1):14–43,

???? 1948. URL http://www.jstor.
org/stable/2980723.

Leslie:1948:DTB

[560] P. H. Leslie. On the distribution in time of the births in successive generations. Journal of the Royal Statistical Society. Series A (General), 111(1): 44–53, ???? 1948. URL http://www.

jstor.org/stable/2980724.

Orcutt:1948:NRA

[561] Guy H. Orcutt. A new regression analyser. Journal of the Royal Statistical Society. Series A (General), 111(1):54–70,

???? 1948. URL http://www.jstor.
org/stable/2980725.

Anonymous:1948:RCNa

[562] Anonymous. Reports from Council and notices to Fellows. Journal of the Royal Statistical Society. Series A (General), 111(1):71–79, ???? 1948. URL http://

www.jstor.org/stable/2980726.

Q:1948:BRB

[563] M. H. Q. Book review: The Fourier Transforms of Probability Distributions, by Aurel Wintner. Journal of the Royal Statistical Society. Series A (General), 111(1):80, ???? 1948. URL http://

www.jstor.org/stable/2980727.

O:1948:BRBa

[564] J. O. Book review: The U.F.A.W.

Handbook on the Care and Management of Laboratory Animals, by Alastair N. Worden; T. Dalling; C. W. Hume. Journal of the Royal Statistical Society. Series A (General), 111(1):80–81, ????

1948. URL http://www.jstor.org/
stable/2980728.

D:1948:BRBa

[565] P. A. D. Book review: Investment, Location, and Size of Plant: A Realistic Inquiry into the Structure of British and American Industries, by P. Sargant Florence; W. Baldamus. Journal of the Royal Statistical Society. Series A (General), 111(1):81, ???? 1948. URL http:

                                        //www.jstor.org/stable/2980729.

D:1948:BRBb

[566] D. J. D. Book review: The Motor Industry of Great Britain, 1947. Journal of the Royal Statistical Society. Series A (General), 111(1):82, ???? 1948.

                                        URL http://www.jstor.org/stable/
                                        2980730.

J:1948:BRB

[567] J. H. J. Book review: Aspects of British Economic History, 1918–1925, by A. C. Pigou. Journal of the Royal Statistical Society. Series A (General), 111(1): 82–84, ???? 1948. URL http://www.

                                        jstor.org/stable/2980731.

Anonymous:1948:SNa

[568] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 111(1):85–87, ???? 1948.

                                        URL http://www.jstor.org/stable/
                                        2980732.

Anonymous:1948:CNa

[569] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 111(1):88–90, ???? 1948.

                                        URL http://www.jstor.org/stable/
                                        2980733.

Anonymous:1948:SEAa

[570] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 111(1):91–94, ???? 1948.

URL http://www.jstor.org/stable/
2980734.

Anonymous:1948:PRa

[571] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 111(1):95–97, ???? 1948.

URL http://www.jstor.org/stable/
2980735.

Anonymous:1948:BMa

[572] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 111(1):??, ???? 1948.

URL http://www.jstor.org/stable/
2980736.

Anonymous:1948:FMb

[573] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 111(2):??, ???? 1948.

URL http://www.jstor.org/stable/
2981022.

Jones:1948:CLS

[574] J. H. Jones. Cost of living subsidies.

Journal of the Royal Statistical Society. Series A (General), 111(2):99–111,

???? 1948. URL http://www.jstor.
org/stable/2981023.

Wallace:1948:FCT

[575] J. Wallace. The fire census and technical field intelligence work of the joint fire research organization. Journal of the Royal Statistical Society. Series A (General), 111(2):112–132, ???? 1948.

URL http://www.jstor.org/stable/
2981024.

Kendall:1948:UKM

[576] M. G. Kendall. United Kingdom merchant shipping statistics. Journal of the Royal Statistical Society. Series A (General), 111(2):133–144, ???? 1948.

                                        URL http://www.jstor.org/stable/
                                        2981025.

Ross:1948:WCC

[577] K. H. Ross. Working class clothing consumption, 1937–1938. Journal of the Royal Statistical Society. Series A (General), 111(2):145–160, ???? 1948.

                                        URL http://www.jstor.org/stable/
                                        2981026.

Anonymous:1948:RCNb

[578] Anonymous. Reports from Council and notices to Fellows. Journal of the Royal Statistical Society. Series A (General), 111(2):161–162, ???? 1948. URL http:

                                        //www.jstor.org/stable/2981027.

C:1948:BRB

[579] E. A. C. Book review: Fundamentals of Statistics, by Truman Lee Kelley. Journal of the Royal Statistical Society. Series A (General), 111(2):163–164,

                                        ???? 1948. URL http://www.jstor.
                                        org/stable/2981028.

O:1948:BRBb

[580] P. D. O. Book review: Principles of Medical Statistics, by A. Bradford Hill. Journal of the Royal Statistical Society. Series A (General), 111(2):164–165,

                                        ???? 1948. URL http://www.jstor.
                                        org/stable/2981029.

K:1948:BRBa

[581] P. H. K. Book review: Europe’s Population in the Inter-war Years, by Dudley Kirk. Journal of the Royal Statistical Society. Series A (General), 111(2):165,

???? 1948. URL http://www.jstor.
org/stable/2981030.

K:1948:BRBb

[582] P. H. K. Book review: Population Policy in Great Britain. Journal of the Royal Statistical Society. Series A (General), 111(2):165–166, ???? 1948. URL http:

//www.jstor.org/stable/2981031.

Anonymous:1948:SNb

[583] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 111(2):167–168, ???? 1948.

URL http://www.jstor.org/stable/
2981032.

Anonymous:1948:CNb

[584] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 111(2):169–170, ???? 1948.

URL http://www.jstor.org/stable/
2981033.

Anonymous:1948:SEAb

[585] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 111(2):171–175, ???? 1948.

URL http://www.jstor.org/stable/
2981034.

Anonymous:1948:PRb

[586] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 111(2):176–179, ???? 1948.

URL http://www.jstor.org/stable/
2981035.

Anonymous:1948:BMb

[587] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 111(2):??, ???? 1948.

                                        URL http://www.jstor.org/stable/
                                        2981036.

Anonymous:1948:FMc

[588] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 111(3):??, ???? 1948.

                                        URL http://www.jstor.org/stable/
                                        2984158.

Anscombe:1948:VCE

[589] F. J. Anscombe. The validity of comparative experiments. Journal of the Royal Statistical Society. Series A (General), 111(3):181–211, ???? 1948. URL http:

                                        //www.jstor.org/stable/2984159.

Wishart:1948:TS

[590] John Wishart. The teaching of statistics. Journal of the Royal Statistical Society. Series A (General), 111(3):212– 229, ???? 1948. URL http://www.

                                        jstor.org/stable/2984160.

Greenwood:1948:SNS

[591] Major Greenwood. The sources and nature of statistical information in special fields of statistics: Medical statistics. Journal of the Royal Statistical Society. Series A (General), 111(3):230– 234, ???? 1948. URL http://www.

                                        jstor.org/stable/2984161.

Nath:1948:PI

[592] Pran Nath. Percentage inspection. Journal of the Royal Statistical Society. Series A (General), 111(3):235–237, ????

                                        1948. URL http://www.jstor.org/
                                        stable/2984162.

Anonymous:1948:PSa

[593] Anonymous. Proceedings of sections. Journal of the Royal Statistical Society. Series A (General), 111(3):238–239,

???? 1948. URL http://www.jstor.
org/stable/2984163.

N:1948:BRBa

[594] J. W. N. Book review: Cours de Statistique Appliquée aux Affaires: V. Statistiques d’Enterprises, by Michel Huber. Journal of the Royal Statistical Society. Series A (General), 111(3):240, ????

1948. URL http://www.jstor.org/
stable/2984164.

L:1948:BRB

[595] H. O. L. Book review: Les Causes de Décès en Suisse, Étudiées a la Lumière de la Démographie Actuelle et de la Démographie Potentielle, by Gérald Mentha. Journal of the Royal Statistical Society. Series A (General), 111(3):240–241, ???? 1948. URL http:

//www.jstor.org/stable/2984165.

B:1948:BRB

[596] B. B. Book review: The Institute of Actuaries 1848–1948, by R. C. Simmonds. Journal of the Royal Statistical Society. Series A (General), 111(3):241–243,

???? 1948. URL http://www.jstor.
org/stable/2984166.

D:1948:BRBc

[597] P. A. D. Book review: Enterprise First, by William Wallace. Journal of the Royal Statistical Society. Series A (General), 111(3):243, ???? 1948. URL http:

//www.jstor.org/stable/2984167.

S:1948:BRBa

[598] G. F. S. Book review: The Theory of Excess Profits Taxation, by M. H. Gopal. Journal of the Royal Statistical Society. Series A (General), 111(3):

243–244, ???? 1948. URL http://www.

jstor.org/stable/2984168.

S:1948:BRBb

[599] G. F. S. Book review: The New Capital Issue Market in India, by M. A. Mulky. Journal of the Royal Statistical Society. Series A (General), 111(3): 244–245, ???? 1948. URL http://www.

jstor.org/stable/2984169.

R:1948:BRB

[600] T. M. R. Book review: Commercial Goodwill: Its History, Value and Treatment in Accounts, by P. D. Leake; Depreciation and Wasting Assets and their Treatment in Computing Annual Profit and Loss. by P. D. Leake. Journal of the Royal Statistical Society. Series A (General), 111(3):245–247, ???? 1948.

URL http://www.jstor.org/stable/
2984170.

Anonymous:1948:SNc

[601] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 111(3):248–250, ???? 1948.

URL http://www.jstor.org/stable/
2984171.

Anonymous:1948:CNc

[602] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 111(3):251, ???? 1948.

URL http://www.jstor.org/stable/
2984172.

M:1948:CEC

[603] P. C. M. Clara Elizabeth Collet. Journal of the Royal Statistical Society. Series A (General), 111(3):252–254, ???? 1948.

URL http://www.jstor.org/stable/
2984173.

Hayek:1948:WCM

[604] F. A. Hayek. Wesley Clair Mitchell, 1874–1948. Journal of the Royal Statistical Society. Series A (General), 111(3): 254–255, ???? 1948. URL http://www.

jstor.org/stable/2984174.

Anonymous:1948:SEAc

[605] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 111(3):256–260, ???? 1948.

URL http://www.jstor.org/stable/
2984175.

Anonymous:1948:VI

[606] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 111(4):393–395, ????

1948. URL http://www.jstor.org/
stable/2980814.

Anonymous:1948:FMd

[607] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 111(4):??, ???? 1948.

URL http://www.jstor.org/stable/
2980815.

Desai:1948:CEI

[608] R. C. Desai. Consumer expenditure in India, 1931–2 to 1940–1. Journal of the Royal Statistical Society. Series A (General), 111(4):261–308, ???? 1948.

URL http://www.jstor.org/stable/
2980816.

Singer:1948:IPE

[609] H. W. Singer and C. E. V. Leser. Industrial productivity in England and Scotland. Journal of the Royal Statistical Society. Series A (General), 111(4):309– 330, ???? 1948. URL http://www.

jstor.org/stable/2980817.

Anonymous:1948:WP

[610] Anonymous. Wholesale prices in 1947. Journal of the Royal Statistical Society. Series A (General), 111(4):331–345,

???? 1948. URL http://www.jstor.
org/stable/2980818.

Anonymous:1948:RCF

[611] Anonymous. Report of the Council for the financial year ended December 31st, 1947, and for the sessional year ending July 20th, 1948, presented at the one hundred and fourteenth Annual General Meeting of the Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine, W.C. 1, on July 20th, 1948. Journal of the Royal Statistical Society. Series A (General), 111 (4):346–363, ???? 1948. URL http://

www.jstor.org/stable/2980819.

Anonymous:1948:POH

[612] Anonymous. Proceedings of the one hundred and fourteenth Annual General Meeting of the Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine on Tuesday, July 20th, 1948. Journal of the Royal Statistical Society. Series A (General), 111 (4):364, ???? 1948. URL http://www.

jstor.org/stable/2980820.

Anonymous:1948:PSb

[613] Anonymous. Proceedings of sections. Journal of the Royal Statistical Society. Series A (General), 111(4):365–368,

???? 1948. URL http://www.jstor.
org/stable/2980821.

F:1948:BRB

[614] D. J. F. Book review: Statistiske Metoder, by A. Hald. Journal of the Royal Statistical Society. Series A (General), 111(4):369, ???? 1948. URL http:

//www.jstor.org/stable/2980822.

N:1948:BRBb

[615] A. M. N. Book review: Planning and the Price Mechanism: The Liberal-Socialist Solution, by James Edward Meade. Journal of the Royal Statistical Society. Series A (General), 111(4): 370–371, ???? 1948. URL http://www.

jstor.org/stable/2980823.

I:1948:BRBa

[616] A. R. I. Book review: Business Finance and Banking, by Neil H. Jacoby; Raymond J. Saulnier. Journal of the Royal Statistical Society. Series A (General), 111(4):371, ???? 1948. URL http://

www.jstor.org/stable/2980824.

I:1948:BRBb

[617] A. R. I. Book review: Studies in Financial Organization, by T. Balogh. Journal of the Royal Statistical Society. Series A (General), 111(4):371–372, ???? 1948.

URL http://www.jstor.org/stable/
2980825.

D:1948:BRBd

[618] P. A. D. Book review: Imperfect Competition in International Trade, by S. B. Rangnekar. Journal of the Royal Statistical Society. Series A (General), 111 (4):372, ???? 1948. URL http://www.

jstor.org/stable/2980826.

I:1948:BRBc

[619] A. R. I. Book review: Municipal Internal Audits, by Arthur Collins. Journal of the Royal Statistical Society. Series A (General), 111(4):372–373, ???? 1948.

                                        URL http://www.jstor.org/stable/
                                        2980827.

Anonymous:1948:ONP

[620] Anonymous. Other new publications. Journal of the Royal Statistical Society. Series A (General), 111(4):373, ????

                                        1948. URL http://www.jstor.org/
                                        stable/2980828.

Anonymous:1948:SNd

[621] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 111(4):374, ???? 1948.

                                        URL http://www.jstor.org/stable/
                                        2980829.

Anonymous:1948:CNd

[622] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 111(4):375, ???? 1948.

                                        URL http://www.jstor.org/stable/
                                        2980830.

G:1948:JWV

[623] R. F. G. James William Verdier, O.B.E. Journal of the Royal Statistical Society. Series A (General), 111(4):376, ????

                                        1948. URL http://www.jstor.org/
                                        stable/2980831.

Anonymous:1948:SEAd

[624] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 111(4):377–382, ???? 1948.

                                        URL http://www.jstor.org/stable/
                                        2980832.

Anonymous:1948:LAL

[625] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 111(4): 383–391, ???? 1948. URL http://www.

jstor.org/stable/2980833.

Anonymous:1948:BMc

[626] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 111(4):??, ???? 1948.

URL http://www.jstor.org/stable/
2980834.

Anonymous:1949:FMa

[627] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 112(1):??, ???? 1949.

URL http://www.jstor.org/stable/
2984176.

Smeed:1949:SSA

[628] R. J. Smeed. Some statistical aspects of road safety research. Journal of the Royal Statistical Society. Series A (General), 112(1):1–34, ???? 1949.

URL http://www.jstor.org/stable/
2984177.

Ainsworth:1949:EWH

[629] R. B. Ainsworth. Earnings and working hours of manual wage-earners in the United Kingdom in October, 1938. Journal of the Royal Statistical Society. Series A (General), 112(1):35–66, ????

1949. URL http://www.jstor.org/

stable/2984178. See erratum [642].

Leak:1949:SNS

[630] H. Leak. The sources and nature of statistical information in special fields of statistics: Statistics of the census of production and distribution. Journal of the Royal Statistical Society. Series A (General), 112(1):67–80, ???? 1949.

URL http://www.jstor.org/stable/
2984179.

Anonymous:1949:PSa

[631] Anonymous. Proceedings of sections. Journal of the Royal Statistical Society. Series A (General), 112(1):81–82,

                                        ???? 1949. URL http://www.jstor.
                                        org/stable/2984180.

A:1949:BRB

[632] P. A. Book review: The Measurement of Colonial National Incomes: An Experiment, by Phyllis Deane. Journal of the Royal Statistical Society. Series A (General), 112(1):83–84, ???? 1949.

                                        URL http://www.jstor.org/stable/
                                        2984181.

B:1949:BRBa

[633] E. J. M. B. Book review: Analysis of Wisconsin Income, by Frank A. Hanna; Joseph A. Pechman; Sidney M. Lerner. Journal of the Royal Statistical Society. Series A (General), 112(1):84–85,

                                        ???? 1949. URL http://www.jstor.
                                        org/stable/2984182.

B:1949:BRBb

[634] E. J. M. B. Book review: Principles of National Income Analysis, by Carl S. Shoup. Journal of the Royal Statistical Society. Series A (General), 112(1): 85–86, ???? 1949. URL http://www.

                                        jstor.org/stable/2984183.

R:1949:BRBa

[635] L. R. Book review: Studies in War Economics. Journal of the Royal Statistical Society. Series A (General), 112(1):86,

                                        ???? 1949. URL http://www.jstor.
                                        org/stable/2984184.

DeN:1949:BRB

[636] A. M. De N. Book review: Towards a Dynamic Economics. Some Recent Developments of Economic Theory and their Application to Policy, by R. F. Harrod. Journal of the Royal Statistical Society. Series A (General), 112(1): 86–88, ???? 1949. URL http://www.

jstor.org/stable/2984185.

Anonymous:1949:SNa

[637] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 112(1):89–91, ???? 1949.

URL http://www.jstor.org/stable/
2984186.

Anonymous:1949:CNa

[638] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 112(1):92–95, ???? 1949.

URL http://www.jstor.org/stable/
2984187.

Anonymous:1949:SEAa

[639] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 112(1):96–100, ???? 1949.

URL http://www.jstor.org/stable/
2984188.

Anonymous:1949:PRa

[640] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 112(1):101–103, ???? 1949.

URL http://www.jstor.org/stable/
2984189.

Anonymous:1949:FMb

[641] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 112(2):??, ???? 1949.

URL http://www.jstor.org/stable/
2981135.

Anonymous:1949:EEW

[642] Anonymous. Erratum: Earnings and working hours of manual wage-earners in the United Kingdom in October, 1938. Journal of the Royal Statistical Society. Series A (General), 112(2):??, ????

                                        1949. URL http://www.jstor.org/
                                        stable/2981136. See [629].

Campion:1949:IS

[643] H. Campion. International statistics. Journal of the Royal Statistical Society. Series A (General), 112(2):105–143,

                                        ???? 1949. URL http://www.jstor.
                                        org/stable/2981137.

Benjamin:1949:LGS

[644] B. Benjamin. Local government statistics–with special reference to the health service. Journal of the Royal Statistical Society. Series A (General), 112 (2):144–179, ???? 1949. URL http://

                                        www.jstor.org/stable/2981138.

Anscombe:1949:TSI

[645] F. J. Anscombe. Tables of sequential inspection schemes to control fraction defective. Journal of the Royal Statistical Society. Series A (General), 112(2): 180–206, ???? 1949. URL http://www.

                                        jstor.org/stable/2981139.

Maizels:1949:SNS

[646] A. Maizels. The sources and nature of statistical information in special fields of statistics: The oversea trade statistics of the United Kingdom. Journal of the Royal Statistical Society. Series A (General), 112(2):207–223, ???? 1949.

                                        URL http://www.jstor.org/stable/
                                        2981140.

Anonymous:1949:PSb

[647] Anonymous. Proceedings of sections.

Journal of the Royal Statistical Society. Series A (General), 112(2):224–225,

???? 1949. URL http://www.jstor.
org/stable/2981141.

W:1949:BRB

[648] D. R. W. Book review: Fundamentals of Statistics, by J. B. Scarborough; R. W. Wagner. Journal of the Royal Statistical Society. Series A (General), 112(2):226,

???? 1949. URL http://www.jstor.
org/stable/2981142.

R:1949:BRBb

[649] W. B. R. Book review: A Statistical Analysis of Advertising Expenditure and the Revenue of the Press, by Nicholas Kaldor; Rodney Silverman. Journal of the Royal Statistical Society. Series A (General), 112(2):227, ???? 1949.

URL http://www.jstor.org/stable/
2981143.

T:1949:BRBa

[650] L. H. C. T. Book review: Comparative Productivity in British and American Industry, by L. Rostas. Journal of the Royal Statistical Society. Series A (General), 112(2):227–228, ???? 1949.

URL http://www.jstor.org/stable/
2981144.

E:1949:BRB

[651] M. J. E. Book review: Demographic Survey of the British Colonial Empire, by R. R. Kuczynski. Journal of the Royal Statistical Society. Series A (General), 112 (2):228, ???? 1949. URL http://www.

jstor.org/stable/2981145.

D:1949:BRB

[652] P. A. D. Book review: Current Financial Problems and the City of London. Journal of the Royal Statistical Society. Series A (General), 112(2):228– 229, ???? 1949. URL http://www.

                                        jstor.org/stable/2981146.

Anonymous:1949:ONP

[653] Anonymous. Other new publications. Journal of the Royal Statistical Society. Series A (General), 112(2):229, ????

                                        1949. URL http://www.jstor.org/
                                        stable/2981147.

Anonymous:1949:SNb

[654] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 112(2):230–231, ???? 1949.

                                        URL http://www.jstor.org/stable/
                                        2981148.

Anonymous:1949:CNb

[655] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 112(2):232–233, ???? 1949.

                                        URL http://www.jstor.org/stable/
                                        2981149.

Anonymous:1949:SEAb

[656] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 112(2):234–238, ???? 1949.

                                        URL http://www.jstor.org/stable/
                                        2981150.

Anonymous:1949:LALa

[657] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 112(2): 239–246, ???? 1949. URL http://www.

                                        jstor.org/stable/2981151.

Anonymous:1949:PRb

[658] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 112(2):247–250, ???? 1949.

URL http://www.jstor.org/stable/
2981152.

Anonymous:1949:FMc

[659] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 112(3):??, ???? 1949.

URL http://www.jstor.org/stable/
2980930.

Tintner:1949:FPS

[660] Gerhard Tintner. Foundations of probability and statistical inference. Journal of the Royal Statistical Society. Series A (General), 112(3):251–286, ???? 1949.

URL http://www.jstor.org/stable/
2980931.

Bronowski:1949:SUS

[661] J. Bronowski. Some uses of statistics in the building industry: An investigation into the erection times of nine types of non-traditional house. Journal of the Royal Statistical Society. Series A (General), 112(3):287–308, ???? 1949.

URL http://www.jstor.org/stable/
2980932.

James:1949:NCT

[662] S. F. James. A note on Carnap’s theory of probability. Journal of the Royal Statistical Society. Series A (General), 112 (3):309–315, ???? 1949. URL http://

www.jstor.org/stable/2980933.

Beales:1949:ICN

[663] R. E. Beales. Industrial classification — national and international. Journal of the Royal Statistical Society. Series A (General), 112(3):316–330, ???? 1949.

URL http://www.jstor.org/stable/
2980934.

George:1949:SNS

[664] R. F. George. The sources and nature of statistical information in special fields of statistics: Statistics relating to the coal mining industry. Journal of the Royal Statistical Society. Series A (General), 112(3):331–337, ???? 1949. URL http:

//www.jstor.org/stable/2980935.

C:1949:BRBa

[665] D. G. C. Book review: Statistics for Economists, by R. G. D. Allen. Journal of the Royal Statistical Society. Series A (General), 112(3):338–339, ???? 1949.

URL http://www.jstor.org/stable/
2980936.

M:1949:BRBa

[666] P. A. M. Book review: Proceedings of the Berkeley Symposium on Mathematical Statistics and Probability. Journal of the Royal Statistical Society. Series A (General), 112(3):339, ???? 1949. URL http:

//www.jstor.org/stable/2980937.

M:1949:BRBb

[667] C. A. M. Book review: Social Surveys, by D. Caradog Jones. Journal of the Royal Statistical Society. Series A (General), 112(3):339–340, ???? 1949.

URL http://www.jstor.org/stable/
2980938.

B:1949:BRBc

[668] W. R. B. and K. F. G. Book review: American Transportation in Prosperity and Depression, by Thor Hultgren. Journal of the Royal Statistical Society. Series A (General), 112(3):340– 342, ???? 1949. URL http://www.

jstor.org/stable/2980939.

C:1949:BRBb

[669] J. R. C. Book review: Expectation in Economics, by G. L. S. Shackle. Journal of the Royal Statistical Society. Series A (General), 112(3):342–343, ???? 1949.

URL http://www.jstor.org/stable/
2980940.

Anonymous:1949:SNc

[670] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 112(3):344–345, ???? 1949.

URL http://www.jstor.org/stable/
2980941.

G:1949:CAV

[671] R. F. G. Conrad Alexander Verrijn Stuart. Journal of the Royal Statistical Society. Series A (General), 112(3):346,

???? 1949. URL http://www.jstor.
org/stable/2980942.

Anonymous:1949:SEAc

[672] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 112(3):347–350, ???? 1949.

URL http://www.jstor.org/stable/
2980943.

Anonymous:1949:LALb

[673] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 112(3): 351–357, ???? 1949. URL http://www.

jstor.org/stable/2980944.

Anonymous:1949:VI

[674] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 112(4):495–496, ????

1949. URL http://www.jstor.org/
stable/2980762.

Anonymous:1949:FMd

[675] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 112(4):??, ???? 1949.

URL http://www.jstor.org/stable/
2980763.

Nicholson:1949:VWC

[676] J. L. Nicholson. Variations in working class family expenditure. Journal of the Royal Statistical Society. Series A (General), 112(4):359–418, ???? 1949.

URL http://www.jstor.org/stable/
2980764.

Kellet:1949:SNS

[677] W. G. G. Kellet. The sources and nature of statistical information in special fields of statistics: International rubber statistics. Journal of the Royal Statistical Society. Series A (General), 112(4): 419–435, ???? 1949. URL http://www.

jstor.org/stable/2980765.

Elderton:1949:FSL

[678] William P. Elderton. A few statistics on the length of English words. Journal of the Royal Statistical Society. Series A (General), 112(4):436–445, ???? 1949.

URL http://www.jstor.org/stable/
2980766.

Moyal:1949:DWT

[679] J. E. Moyal. The distribution of wars in time. Journal of the Royal Statistical Society. Series A (General), 112(4): 446–449, ???? 1949. URL http://www.

jstor.org/stable/2980767.

Anonymous:1949:WP

[680] Anonymous. Wholesale prices in 1948.

Journal of the Royal Statistical Society. Series A (General), 112(4):450–464,

???? 1949. URL http://www.jstor.
org/stable/2980768.

Anonymous:1949:RC

[681] Anonymous. Report of the Council.

Journal of the Royal Statistical Society. Series A (General), 112(4):465–481,

???? 1949. URL http://www.jstor.
org/stable/2980769.

Anonymous:1949:POH

[682] Anonymous. Proceedings of the one hundred and fifteenth Annual General Meeting of the Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine, on Wednesday, June 29th, 1949. Journal of the Royal Statistical Society. Series A (General), 112(4):482, ???? 1949. URL http://

www.jstor.org/stable/2980770.

T:1949:BRBb

[683] L. H. C. T. Book review: Sampling Methods for Censuses and Surveys, by F. Yates. Journal of the Royal Statistical Society. Series A (General), 112(4): 483–484, ???? 1949. URL http://www.

jstor.org/stable/2980771.

F:1949:BRB

[684] F. A. F. Book review: Statistics and their Application to Commerce, by A. Lester Boddington. Journal of the Royal Statistical Society. Series A (General), 112(4):484, ???? 1949. URL http://

www.jstor.org/stable/2980772.

N:1949:BRB

[685] J. W. N. Book review: Die Geschichte der Lage der Arbeiter in Deutschland von 1800 bis in die Gegenwart, by J. Kuczynski. Journal of the Royal Statistical Society. Series A (General), 112 (4):485, ???? 1949. URL http://www.

                                        jstor.org/stable/2980773.

Anonymous:1949:CNc

[686] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 112(4):486, ???? 1949.

                                        URL http://www.jstor.org/stable/
                                        2980774.

H:1949:MG

[687] A. B. H. and William Butler. Major Greenwood. Journal of the Royal Statistical Society. Series A (General), 112 (4):487–489, ???? 1949. URL http://

                                        www.jstor.org/stable/2980775.

Anonymous:1949:SEAd

[688] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 112(4):490–494, ???? 1949.

                                        URL http://www.jstor.org/stable/
                                        2980776.

Anonymous:1949:BM

[689] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 112(4):??, ???? 1949.

                                        URL http://www.jstor.org/stable/
                                        2980777.